### v5

# Semi-supervised Underwater Image Enhancement Using A Physics-Aware Triple-Stream Network

Abstract
Underwater images normally suffer from degradation due to the transmission medium of water bodies. Both traditional prior-based approaches and deep learning-based methods have been used to address this problem. However, the inflexible assumption of the former often impairs their effectiveness in handling diverse underwater scenes, while the generalization of the latter to unseen images is usually weakened by insufficient data. In this study, we leverage both the physics-based Image Formation Model (IFM) and deep learning techniques for Underwater Image Enhancement (UIE). To this end, we propose a novel Physics-Aware Triple-Stream Underwater Image Enhancement Network, i.e., PATS-UIENet, which comprises a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Light Estimation Stream (A-Stream). This network fulfills the UIE task by explicitly estimating the degradation parameters of a revised IFM. We also adopt an IFM-inspired semi-supervised learning framework, which exploits both the labeled and unlabeled images, to address the issue of insufficient data. To our knowledge, such a physics-aware deep network and the IFM-inspired semi-supervised learning framework have not been used for the UIE task before. Our method performs better than, or at least comparably to, sixteen baselines across six testing sets in the degradation estimation and UIE tasks. These promising results should be due to the fact that the proposed method can not only model the degradation but also learn the characteristics of diverse underwater scenes. [^1] The models and code will be published on the acceptance of the paper.
Index Terms:
Underwater Image Enhancement (UIE), Underwater Image Processing, Image Formation Model (IFM), Deep learning, Semi-supervised Learning.
I Introduction
The images captured in the underwater environment play important roles in ocean exploration. However, these images normally suffer from different degradation, due to the wavelength-dependent light absorption and light scattering caused by the underwater transmission medium. According to the Image Formation Model (IFM) [1], the process of image degradation can be formulated as follows:
$I^{c}(x)=J^{c}(x)t^{c}(x)+(1-t^{c}(x))A^{c},$ (1)
where $I^{c}(x)$ ($c\in\{R,G,B\}$) is a degraded image, $J^{c}(x)$ is the underlying clean image, $t^{c}(x)=exp(-\beta^{c}d(x))$ denotes the transmission maps, $\beta^{c}$ is the attenuation factor for each channel, $d(x)$ is the scene distance map and $A^{c}$ is the ambient light. (See Fig. 1 for examples). In this context, $J^{c}(x)t^{c}(x)$ describes the distance-dependent color distortion, while $(1-t^{c}(x))A^{c}$ represents the scattering of the ambient light which decreases the contrast and lessens the visibility of the image. Despite the IFM had been widely used, its flaws in accurately modeling the degradation process was highlighted [2].
Figure 1: Examples of a degraded underwater image, the corresponding ambient light image $A^{c}$ and three transmission maps $t^{c}(x)$ ($c\in\{R,G,B\}$) estimated, and the enhanced image produced using the estimated parameters of the IFM [1].
Considering that the degradation may interfere with downstream tasks, the design of effective UIE methods is critical yet challenging. Existing methods can be divided into two categories, i.e., traditional prior-based methods and deep learning-based methods. Traditional prior-based methods [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] typically rely on a certain form of prior information. Some prior-based methods [3, 4, 5] aim at estimating the parameters of the IFM. However, they normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene violates the prior assumption [5]. On the other hand, the other prior-based methods [11, 12, 13, 14, 15] directly enhance the quality of underwater images without explicitly modeling the degradation process. Due to the lack of the knowledge of underwater scenes and the limited adaptability and robustness, they cannot effectively enhance the images captured in the complex scenes and often introduce artifacts.
In contrast, deep learning-based UIE approaches [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] aim to directly learn from different types of degraded images. Therefore, the problem with the limited static priors can be alleviated. Ideally, a large number of real-world labeled training images should be utilized. However, it is difficult to collect such a large labeled underwater data set in practice. Alternatively, the best enhanced image can be selected as the reference (ground-truth) by observers from the images processed by a set of UIE algorithms [22, 31]. Nevertheless, this sort of data sets cannot represent diverse underwater scenes due to the limited size. In this case, the deep UIE model trained may overfit the training data.
The synthetic images generated using the IFM [32, 18] or a Generative Adversarial Network (GAN) [33, 34, 35, 20] were used to address the problem with the lack of labeled data. However, the utilization of these images to simulate real-world underwater degradation processes is struggling. Transfer learning techniques [34, 19] were also applied for that purpose, which usually used generative models to map degraded images to the clearer images. But they are difficult to train and may produce unstable results. Although unsupervised learning methods [36, 37] alleviate the reliance on the paired data by directly learning degradation parameters from unlabeled degraded images, they often suffer from slow inference speed and are prone to color distortion. Recently, semi-supervised learning methods [38, 28, 29, 30], which combine the strengths of both the supervised and unsupervised approaches, have achieved promising results by leveraging the labeled data along with the unlabeled data during the training process.
Despite the progress has been made by the existing methods, two challenges remain. First, existing deep learning-based methods lack effective physics-aware modeling, which limits their interpretability and effectiveness in modeling the degradation process. Second, a supervised or unsupervised approach normally struggles with the insufficient or noisy data due to the complexity of real underwater degradation, leading to overfitting or weak generalization.
To address these challenges, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network, i.e., PATS-UIENet, which explicitly incorporates the revised IFM [2] into the deep neural network that we deliberately design. Unlike existing physics-based methods [3, 4, 5], our network explicitly estimates the parameters of a physical model using three streams, including a Direct Signal Transmission Estimation Stream (D-Stream), a Backscatter Signal Transmission Estimation Stream (B-Stream) and an Ambient Estimation Stream (A-Stream). Due to the explicit modeling of the physical degradation process, our network is superior to existing UIE methods by exploiting the strengths of both the theoretical interpretability of physics-based approaches and the powerful feature representation ability of deep neural networks.
We further introduce an IFM-inspired semi-supervised learning framework, including a bi-directional supervised scheme and an unsupervised scheme, for the purpose of overcoming the limitation of insufficient data. The supervised scheme effectively utilizes the limited labeled data by providing explicit degradation guidance, while the unsupervised scheme exploits the abundant unlabeled data to improve generalization. Therefore, our framework can be better trained than the existing supervised approaches [18, 19, 22, 23, 24, 25, 26, 27], unsupervised methods [36, 37] and semi-supervised approaches [38, 28, 29, 30] using the same amount of data. As a result, the generalization of the model trained using our framework is stronger than that trained using existing supervised, unsupervised and semi-supervised approaches.
To our knowledge, this study makes the first effort to jointly apply the physics-aware deep network and the IFM-inspired semi-supervised learning technique to the UIE task. Our main contributions can be summarized as threefold.
  • We introduce a Physics-Aware Triple-Stream Underwater Image Enhancement Network, referred to as PATS-UIENet, which can explicitly estimate the degradation parameters of the revised IFM. This network differentiates itself from existing IFM-based methods by integrating physical parameter estimation into a deliberately designed deep neural network. Such an IFM-motivated network has not been explored for the UIE task before.
  • We propose an IFM-inspired semi-supervised learning framework, which addresses the issues of data insufficiency and training instability by exploiting the merits of both the supervised and unsupervised learning methods.
  • We conduct a series of comparative experiments on six underwater testing sets along with sixteen baselines. The results not only validate the effectiveness of our method but also provide benchmarks for future research.
The remainder of this paper is organized as follows. The related literature is reviewed in Section II. We introduce the proposed method in Section III. The experimental settings and results are reported in Sections IV and V respectively. Finally, our conclusion is drawn in Section VI.
II Related Work
II-A Revised Image Formation Model
As pointed out by Akkaynak and Treibitz [2], the original IFM [1] ignored the dependencies of the backscatter coefficient on the ambient light and the optical property of water bodies. They assumed that the attenuation coefficients of the direct signal and the backscattered signal are the same. In this case, the original IFM [1] cannot adequately describe the degradation process of underwater images. It has been demonstrated that the attenuation factors of the direct signal and the backscatter signal are different while the attenuation factor of the backscatter signal is affected more severely by the ambient light [2].
Akkaynak and Treibitz [2] further proposed a revised IFM, which can be expressed as:
$I^{c}(x)=J^{c}(x)e^{-\beta_{c}^{D}(\mathbf{D})\cdot z}+\left(1-e^{-\beta_{c}^{B}(\mathbf{B})\cdot z}\right)A^{c},$ (2)
where $\textbf{D}=\{z,\rho,E,S_{c},\beta\}$ and $\textbf{B}=\{E,S_{c},b,\beta\}$ are two sets of parameters which affect the attenuation factors of the direct signal $\beta_{c}^{D}$ and the backscatter signal $\beta_{c}^{B}$, respectively, $z$ represents the distance between the scene and the camera, $\rho$ is the reflectance spectrum of the object, $E$ is the ambient light at a certain distance, $S_{c}$ is the camera response function, and $b$ and $\beta$ are the beam scattering and attenuation coefficients, respectively.
In contrast to the original IFM [1], the revised IFM [2] offers the more comprehensive and physically accurate representation of underwater image degradation. Therefore, we design the proposed method on top of the revised IFM. For more details, please refer to the original publication [2].
II-B Prior-Based Methods
To recover clear underwater images, many methods [3, 4, 5, 6, 7, 8, 10] used the prior information to estimate the degradation parameters of the IFM. Although the Dark Channel Prior (DCP) [3] was initially proposed for image dehazing, some researchers [6, 7] employed it for underwater image enhancement, due to the similarity between the degradation processes resulted from the foggy weather and underwater environment. However, the original DCP usually yielded erroneous estimations. Therefore, more studies were performed to improve it, such as Underwater Dark Channel Prior (UDCP) [4] and Generalization of the Dark Channel Prior (GDCP) [5]. Moreover, the Haze Line Prior (HL) [8] was utilized in some studies [10, 39]. Despite these methods were designed on top of the IFM, they normally made a rigid assumption about the underwater environment, which restricted the application of them to a specific underwater scenario.
On the other hand, some approaches used the more general priors [40, 11, 12, 13, 16, 14, 15] to enhance the quality of underwater images without taking the IFM into account, including Histogram Equalization (HE) [40, 11], Retinex-based methods [13, 16] and image fusion techniques [17, 41]. These approaches usually enhanced the contrast and produced more color-balanced results. However, they probably struggle with processing globally inhomogeneous degraded images and may introduce artifacts, such as halos and color casts, due to the lack of the knowledge of underwater scenes.
II-C Learning-Based Methods
Thanks to the powerful representation learning ability and large-scale training data [42], deep learning-based methods greatly boosted the development of computer vision in both high-level [43, 44, 45] and low-level vision tasks [46, 47, 48]. However, the application of deep learning to underwater image enhancement is much less than other tasks. This dilemma should be attributed to the difficulty on collecting a large number of underwater images and the corresponding clean counterparts. Some deep learning methods [32, 49] used the fake underwater images, synthesized based on terrestrial images, to train a network for degradation parameter estimation. Since the synthesis algorithms were relatively simple, they usually could not simulate diverse underwater scenes. Besides, the content of the terrestrial images was significantly different from underwater scenes. As a result, those methods usually encountered the domain-shift problem when they were applied to real underwater images.
To overcome this problem, some labeled real-world underwater image data sets, e.g., UIEB [22] and SUIM-E [31], were collected. Regarding a degraded image, a set of enhancement methods were applied. The most visually pleasant result was manually picked out as the reference of the degraded image. A UIE model can be trained using these data sets without considering the IFM. For example, the WaterNet [22] used CNNs to learn fusion weights for the purpose of fusing the enhanced results produced by the White Balance [17], HE and Gamma Correction techniques. To perform the UIE task, the UColor [23] method combined the RGB, HSV and Lab color spaces with the guidance of the IFM parameters estimated using the GDCP [5]. Since the training data sets were relatively small, these methods usually overfitted the training data and poorly generalized to unseen images. In addition, they could not explicitly utilize the information provided by the IFM.
In contrast to the lack of labeled underwater images, it is easier to collect unlabeled underwater images. Existing studies [19, 50] sourced real-world underwater images and categorized them according to the degree of degradation. Unsupervised learning methods [24, 36] were developed on top of these data by learning the mapping from severely degraded images to slightly degraded images. However, these approaches were difficult to train and tended to produce unstable results. Recently, semi-supervised learning methods [38, 28, 29, 30] have been developed, using both labeled and unlabeled underwater images. Although promising results were derived, these methods did not consider physical principles.
The above-mentioned studies either are not robust to diverse underwater scenes, or lack the knowledge of underwater images. We are hence motivated to exploit both the IFM-inspired semi-supervised learning technique and the physics-aware deep neural network, to address these issues.
III Methodology
Considering the importance of physical principles to the UIE task, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network (PATS-UIENet). This network contains a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Light Estimation Stream (A-Stream), which are used to explicitly estimate the three parameters of the revised IFM [2], respectively. To address the challenge of the lack of labeled real-world underwater images, we further adopt an IFM-inspired semi-supervised learning framework, which consists of a bi-directional supervised scheme and an unsupervised scheme. Compared to the supervised or unsupervised method, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model trained is thus stronger, due to the complementary action of the two schemes.
III-A Physics-Inspired Design
Motivated by the revised IFM [2], we propose a physics-aware underwater image enhancement network on top of this model. Since the network explicitly estimates the parameters of the revised IFM by learning from training images, it avoids the rigid constraints of priors while improving the interpretability of the model trained. To reduce the complexity of the network, we simplify Eq. (2) as follows:
$I^{c}(x)=J^{c}(x)t_{D}^{c}(x,\textbf{D})+\left(1-t_{B}^{c}(x,\textbf{B})\right)A^{c}.$ (3)
where $I^{c}(x)$ ($c\in\{R,G,B\}$) is the degraded image, $J^{c}(x)$ is the underlying clean image, $t_{D}^{c}(x,\textbf{D})$ and $t_{B}^{c}(x,\textbf{B})$ represent the direct and backscatter signal transmission maps, respectively, and $A^{c}$ denotes the ambient light. To estimate the three parameters, including $t_{D}^{c}(x,\textbf{D})$, $t_{B}^{c}(x,\textbf{B})$ and $A^{c}$, we design a D-Stream, a B-Stream and an A-Stream, respectively.
We also introduce an IFM-inspired semi-supervised learning framework by leveraging the explicitly estimated parameters of the revised IFM [2]. This framework consists of a bi-directional supervised learning scheme and an unsupervised learning scheme, which addresses the challenge of rare labeled real-world underwater images. Compared to the supervised or unsupervised methods, our PATS-UIENet can take advantage of both labeled and unlabeled real-world images for the more effective training operation. Furthermore, the complementary action between the two schemes enhances the generalization ability of the model trained.
III-B PATS-UIENet
As illustrated in Fig. 2, the PATS-UIENet contains two encoder-decoder style streams with the same structure, i.e., D-stream and B-stream, and a Transformer-based A-stream. The three streams are used to estimate the direct signal transmission map, the backscatter signal transmission map and the ambient light, respectively. Since the red channel usually encounters more severe attenuation than the blue and green channels [1, 51], the utilization of this channel is key to color restoration. Therefore, a simple Red Channel Tuner (RCT) module is designed to adaptively emphasize the red channel before an image is fed into the transmission estimation steams.
The output of the first encoder block $Enc\_1$ in the B-stream is processed by Patchify and the result is sent to the A-stream. A Residual Communication Module (RCM) connects the two blocks of the encoder of the B-stream and the A-stream at the same level, to fulfill the bi-directional feature exchange. The two sets of transmission maps estimated using the D-stream and B-stream, respectively, and the ambient light estimated using the A-stream are fed into the revised IFM [2]. As a result, the enhanced underwater image is produced.
Figure 2: The architecture of the proposed PATS-UIENet, which comprises three individual streams, namely, D-Stream, B-Stream and A-Stream, to estimate the degradation parameters of the revised IFM [2].
Red Channel Tuner (RCT). Inspired by the channel attention mechanism [52], we propose a compact RCT module in order to adaptively emphasize the information contained in the red channel. This module first uses a convolutional layer to extract basic features from the input image, and then use the Global Average Pooling (GAP) to obtain a global representation. A fully-connected layer and the Sigmoid activation function are further used to transform this representation to a tuning weight. The weight is finally used to scale the red channel of the image.
D-Stream. D-stream is designed to estimate the direct signal transmission map of an underwater image. According to Eq. (3), the transmission map varies at the locations of different pixels. In this case, the network should be able to generate fine-grained representations. Hence, we build the D-Stream using a CNN-based encoder-decoder network, due to its strong ability to learn local characteristics, which helps preserve edges and fine details. The encoder comprises five consecutive blocks, denoted as $Enc\_1$ to $Enc\_5$. As a result, multi-scale feature maps are derived using the encoder. Given $Enc\_5$ is used as the bottleneck, the decoder contains four blocks, denoted as $Dec\_1$ to $Dec\_4$, symmetrically. Skip connection is used to pass the feature maps produced by an encoder block to the corresponding decoder block at the same level, which is useful for restoring the fine-grained spatial structure. The use of multi-scale features and skip connections enhances the ability of the network to restore the spatial structure of the scene. The output of the last decoder block is fed into a convolutional layer. Three transmission maps are produced in terms of different color channels. In essence, D-Stream is specifically focused on modeling the directly transmitted light, which normally contains the spatial structure of the scene.
B-Stream. B-stream is responsible for modeling the degradation caused by the backscattering in underwater environments. Although it shares the same CNN-based architecture as D-Stream for training stability, B-stream differs in functionality and processing strategy. Unlike direct signals, backscatter signals exhibit strong global interference and require a context-aware estimation. To this end, we pass the output of the first encoder block $Enc\_1$ through a Patchify processing and feed it into the A-Stream, leveraging the global modeling capability of Transformer to enhance the estimation accuracy. Moreover, B-Stream is connected with the A-Stream via Residual Communication Modules (RCMs), enabling bidirectional feature exchange. Compared to the D-Stream, B-Stream is focused on modeling the global degradation rather than recovering the local scene content.
A-Stream. Ambient light reflects the overall luminance in the underwater scene and is generally spatially homogeneous. To effectively model this global characteristic, we employ a Transformer-based design instead of CNNs, because Transformer [53] is well-suited for capturing long-range dependencies and global characteristics through the self-attention mechanism. A-Stream consists of five Transformer blocks, denoted as $Trans\_1$ to $Trans\_5$, and includes a dedicated “Ambient” token to specifically learn the ambient light parameter. Additionally, A-Stream is coupled with the B-Stream through the RCM, allowing it to incorporate local features and mitigate the lack of spatial detail inherent in Transformer. A-Stream is distinguished from the other streams by its global feature learning process and Transformer-based architecture, which complements the local characteristic modeling of the CNN-based streams.
Residual Communication Module (RCM). According to the revised IFM [2], the degradation can be considered as the outcome caused by both the transmission and the ambient light. In this situation, the backscatter signal transmission estimation steam and the ambient light estimation stream will learn some common features. Therefore, the feature exchange between both the streams is likely to boost the training of each stream. Recently, the complementary action of CNNs and Transformers has been shown [54]. We are inspired to propose the RCM for the sake of bridging the two streams. Specifically, the tokens produced by a block in the A-stream are folded into a set of feature maps while the feature maps generated by a block of the encoder of the B-stream are downsampled to the shape of these maps. Then both sets of maps are concatenated along the channel dimension and are fed into a convolutional layer. Finally, the resultant maps are split into two sets along the channel dimension. Each set is added with the original feature maps of the related stream.
III-C IFM-Inspired Semi-supervised Learning Framework
To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig. 3), inspired by the revised IFM [2] theory, embedding physical constraints directly into the learning process to effectively leverage unlabeled data. This framework includes a bi-directional supervised scheme and an unsupervised scheme. The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated parameters of the revised IFM to perform UIE, and a backward-degradation module, which uses the estimated parameters to degrade the reference images to their real-world degraded counterparts. As a result, the bi-directional supervised scheme can learn a more accurate estimation of the parameters using limited real-world data than that learned using a supervised method with a large number of synthetic data. Also, the generalization of the model trained is stronger than that trained using a supervised method with limited real-world data. On the other hand, the unsupervised scheme exploits the unlabeled real-world underwater images that we collect. A second set of images with different degrees of degradation are generated from these images based on the revised IFM [2]. Two sets of parameters can be estimated using both sets of images, respectively. Each set is utilized as the reference for the other set because their IFMs are related.
Figure 3: The proposed semi-supervised learning framework, which comprises a bi-directional supervised learning scheme and an unsupervised learning scheme, used for training our PATS-UIENet.
Bi-directional Supervised Learning. Given a set of labeled training data, we first use the estimated parameters to obtain enhanced results $\hat{J}^{c}(x)$. Then we perform the forward-enhancement learning module by minimizing the following loss function:
$\mathcal{L}_{fwd}=||(J^{c}(x)-\hat{J}^{c}(x))||_{2}^{2}.$ (4)
Due to the ill-posed nature of the revised IFM, however, only $\mathcal{L}_{fwd}$ may be insufficient for ensuring the validity of the estimated parameters. Therefore, we propose a backward-degradation learning module which degrades clear reference images using the estimated parameters to their degraded counterparts $\hat{I}^{c}(x)$. The supervision over the backward-degradation is achieved by minimizing the loss function:
$\mathcal{L}_{bwd}=||(I^{c}(x)-\hat{I}^{c}(x))||_{2}^{2}.$ (5)
Inspired by the Retinex [13] theory, we further apply Gaussian blur to input images in order to derive the hint of the global ambient light, which is able to boost the training of the A-stream. Given the ambient light $\hat{A}^{c}$ estimated using the A-stream, we minimize the following loss function:
$\mathcal{L}_{A\text{-}sup}=||(I^{c}(x)*G-\hat{A}^{c}||_{2}^{2},$ (6)
where $G$ stands for a Gaussian kernel and * is the convolution operation. Finally, the bi-directional supervised learning scheme is conducted as:
$\mathcal{L}_{sup}=\mathcal{L}_{fwd}+\lambda_{1}\mathcal{L}_{bwd}+\lambda_{2}\mathcal{L}_{A\text{-}sup},$ (7)
where $\lambda_{1}$ and $\lambda_{2}$ are used to balance different loss functions. Compared with the supervised methods [32, 22, 23], this scheme exploits the limited labeled underwater images better.
Unsupervised Learning. To leverage unlabeled underwater images, we adopt an unsupervised learning scheme based on the physical modeling capability of the revised IFM [2] rather than using generative networks [34, 19], which are difficult to train and often produce unstable results. This scheme adaptively constructs pairs of images with different levels of degradation from the unlabeled data.
Given an unlabeled real-world degraded image $I^{c}_{1}(x)$, the revised IFM can be expressed as:
$I^{c}_{1}(x)=J^{c}_{1}(x)t_{D1}^{c}+(1-t_{B1}^{c}(x))A^{c}_{1}.$ (8)
A new degraded image $I^{c}_{2}(x)$ can be purposely derived according to:
$I^{c}_{2}(x)=\alpha I^{c}_{1}(x)+(1-\alpha)(1-t_{B1}^{c}(x))A^{c}_{1},$ (9)
where $\alpha\in(0,1)$ is a controlled factor which decides the degradation extent. Substituting $I^{c}_{1}(x)$ in Eq. (8) into Eq. (9), $I^{c}_{2}(x)$ can be then expressed as:
$I^{c}_{2}(x)=J^{c}_{1}(x)(\alpha t_{D1}^{c}(x))+(1-t_{B1}^{c}(x))A^{c}_{1}.$ (10)
Eq. (10) also satisfies the revised IFM but presents a more severe degradation than the $I^{c}_{1}(x)$, generated from the same underlying clear image $J^{c}_{1}(x)$, because $\alpha t_{D1}^{c}(x)<t_{D1}^{c}(x)$ always holds true. Therefore, the parameters $\hat{t}_{D2}^{c}(x)$ and $\hat{t}_{B2}^{c}(x)$ estimated using the PATS-UIENet from image $I^{c}_{2}(x)$ will be close to $\alpha\hat{t}_{D1}^{c}(x)$ and $\hat{t}_{B1}^{c}(x)$, respectively, where $\hat{t}_{D1}^{c}(x)$ and $\hat{t}_{B1}^{c}(x)$ are the parameters estimated from $I^{c}_{1}(x)$. Correspondingly, two loss functions are defined as:
$\mathcal{L}_{D}=||\hat{t}_{D2}^{c}(x)-\alpha\hat{t}_{D1}^{c}(x)||_{2}^{2},$ (11)
$\mathcal{L}_{B}=||\hat{t}_{B1}^{c}(x)-\hat{t}_{B2}^{c}(x)||_{2}^{2}.$ (12)
Similar to Eq. (6), the loss function for the ambient light $\hat{A}^{c}$ estimated by the A-stream is defined as:
$\mathcal{L}_{A\text{-}unsup}=||(I^{c}(x)*G-\hat{A}^{c}||_{2}^{2}.$ (13)
Due to the lack of reference images, we use a non-reference loss function [37] based on the Gray-world prior [55] to constrain the enhancement of image $\hat{J}^{c}_{1}(x)$, which is expressed as:
$\mathcal{L}_{gw}=\displaystyle\sum_{c\in\{R,G,B\}}||E(J^{c}_{1}(x))-0.5||_{2}^{2},$ (14)
where $E(J^{c}_{1}(x))$ denotes the average of a specific channel $c$. Finally, we define the loss function of the unsupervised learning scheme as:
$\mathcal{L}_{unsup}=\mathcal{L}_{D}+\mathcal{L}_{B}+\mathcal{L}_{A\text{-}unsup}+\lambda_{3}\mathcal{L}_{gw},$ (15)
where $\lambda_{3}$ is a weighting factor.
Semi-supervised Learning. On top of both the bi-directional supervised learning and unsupervised learning schemes, the proposed semi-supervised learning framework can be formulated as:
$\mathcal{L}_{semi\text{-}sup}=\mathcal{L}_{sup}+\lambda_{unsup}\mathcal{L}_{unsup},$ (16)
where $\lambda_{unsup}$ is used to balance the two schemes.
IV Experimental Settings
In this section, we will briefly introduce the baselines, data sets, evaluation metrics and implementation details utilized in our experiments.
IV-A Baselines
We compared the proposed method with two prior-based approaches which were developed on top of the IFM [1], including UDCP [4] and DHL [10], three prior-based methods which did not consider the IFM, including Histogram Equalization (HE) [9], HLRP [15] and MMLE [14]. We also compared our method with seven supervised learning approaches, including WaterNet [22], UColor [23], FUnIE-GAN [20], PUIE-Net (MP) [25], U-Transformer [21], URanker [26] and CCMSRNet [27], an unsupervised learning method, i.e., USUIR [24], and three semi-supervised learning methods, including DDFormer [29], Semi-UIR [28] and UWFormer [30].
IV-B Data Sets
We conducted a series of experiments using five data sets, including four publicly available real-world underwater data sets, i.e., UIEB [22], SUIM-E [31], RUIE [56] and SQUID [10], and a synthetic data set, namely, EUVP [20]. The UIEB data set originally contains 890 pairs of degraded images, in which each reference image was selected from the results produced by applying 12 algorithms to a degraded image. Ten pairs of UIEB [22] images were discarded because the degraded images in these pairs had been included in the rest pairs. In total, we only used 880 pairs of UIEB images.
The SUIM-E [31] data set comprises 1,525 pairs of degraded underwater images and a testing set of 110 labeled images. Three subsets are comprised of the RUIE [56] data set, including UIQS, UCCS and UHTS, which include 3,630, 300 and 300 unlabeled images, respectively. The SQUID [10] data set comprises 114 underwater images captured using a stereo camera, divided into the Michmoret, Katzaa, Nachsholim and Satil subsets. We utilized 2,185 images from the Underwater Scenes subset of the EUVP [20] data set.
Two different PATS-UIENet models were trained for the real-world and synthetic testing images, referred to as $PATS\text{-}UIENet_{Real}$ and $PATS\text{-}UIENet_{Syn}$, respectively. Regarding both the models, we randomly selected 720 pairs from the 880 pairs of UIEB [22] images and 2,000 pairs from the Underwater Scenes subset of the EUVP [20] data set, respectively, which were used as the training images in the supervised learning scheme. We collected 1,934 unlabeled real-world underwater images, which cover diverse underwater scenes, and used these images in the unsupervised learning scheme for both the models.
Among the rest of the UIEB [22] data set, 80 pairs of images were randomly selected as the validation set, while the remaining pairs were used as the testing set which is referred to as Test-U80. The UIEB data set also consists of 60 challenging images without references, which were used as the second testing set, referred to as Test-C60. A third testing set was obtained from the testing set of the SUIM-E data set [31], which contains 110 pairs of labeled images and is named as Test-S110. For the RUIE [56] data set, the UCCS was utilized as the fourth testing set, denoted as Test-UCCS. Fifty-three SQUID [10] images captured using the camera at the right hand side were randomly selected and were used as the fifth testing set, namely, Test-R53. The remaining 185 images in the Underwater Scenes subset of the EUVP [20] data set were used as the sixth testing set, referred to as Test-Scenes.
IV-C Evaluation Metrics
For the labeled testing data, we used Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [57] and Learned Perceptual Image Patch Similarity (LPIPS) [58] to assess the quality of an enhanced image with regard to the reference image. Regarding the unlabeled testing data, we used the Underwater Image Fidelity (UIF) [59] and Multi-scale Image Quality (MUSIQ) [60] metrics. When the SQUID [10] data set was tested, the Average Reproduction Angular Error ($\bar{\psi}$) [10] was used to evaluate the quality of color restoration. Since transmission maps are related to the scene distance, we used the Pearson Correlation Coefficient (PCC) calculated between the estimated transmission maps and the ground-truth depth map to assess the performance of transmission estimation [10].
IV-D Implementation Details
We implemented the proposed method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3090 graphics processing unit. During the training process, the images were resized to a resolution of $256\times 256$ pixels. The number of filters in the RCT was set to 16. For the $Enc\_1$ to $Dec\_1$ in the D-Stream and B-Stream, the numbers of filters were set to 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn. Regarding the A-Stream, the dimension of each token was 384 and the number of attention heads in the MHSA was 6. We trained the PATS-UIENet using the AdamW [61] optimizer. The learning rate and batch size were set to 1e-4 and 12, respectively. Four weighting factors, including $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{unsup}$, were set to 0.1, 0.005, 1 and 0.1, respectively. We first trained the PATS-UIENet using the bi-directional supervised scheme for 50 epochs as a warm-up stage. Then the semi-supervised scheme was used to train it for 1000 epochs. When the semi-supervised learning process was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for each epoch.
V Experimental Results
In this section, we will report the results obtained in the underwater image enhancement, transmission estimation and color restoration experiments and the ablation study.
V-A Underwater Image Enhancement
We evaluated the proposed PATS-UIENet along with sixteen baselines using three full-reference quantitative metrics and two non-reference quantitative metrics. A qualitative analysis and a performance analysis were also performed. In this subsection, the results obtained in the four experiments will be reported.
V-A1 Full-Reference Quantitative Evaluation
Since the Test-U80, Test-S110 and Test-Scenes testing sets contain reference images, a full-reference quantitative evaluation was conducted on the proposed method and the 16 baselines using the PSNR, SSIM [57] and LPIPS [58] metrics across these testing sets. The results are shown in Table I. As can be seen, our PATS-UIENet normally achieved the best performance across all three metrics, regardless of the testing set. Compared with the baselines designed on top of the IFM [1], such as UDCP [4] and DHL [8], and the existing semi-supervised learning approaches, including DDFormer [29], Semi-UIR [28] and UWFormer [30], our method normally showed a large margin.
TABLE I: The full-reference quantitative evaluation of the proposed PATS-UIENet and sixteen baselines on three real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts, respectively.
Method Test-U80 Test-S110 Test-Scenes
[…]
UDCP [4] 9.51 33.66 41.74 10.07 34.29 37.66 14.76 56.43 30.31
DHL [10] 15.16 63.93 30.40 14.61 62.93 28.39 14.99 69.73 22.36
HE [9] 16.60 77.86 25.67 15.41 74.79 30.05 13.61 62.82 36.10
HLRP [15] 13.56 22.76 33.90 12.55 29.24 33.18 12.12 18.21 39.35
MMLE [14] 18.56 76.21 22.57 17.32 77.39 22.77 14.89 62.23 31.99
WaterNet [22] 17.13 70.41 40.01 18.83 75.22 31.08 25.84 84.02 16.28
UColor [23] 21.05 85.24 17.97 20.26 84.25 15.07 24.31 81.40 19.05
FUnIE-GAN [20] 15.31 59.68 44.28 16.89 66.32 34.98 25.11 83.16 13.76
USUIR [24] 16.92 68.69 28.63 17.54 76.58 21.38 15.14 66.51 32.44
PUIE-Net (MP) [25] 22.26 88.89 12.01 21.85 89.19 9.70 21.85 73.47 24.45
U-Transformer [21] 20.98 72.76 26.32 19.92 68.25 30.17 24.09 79.98 15.08
URanker [26] 21.87 85.75 18.64 21.27 87.43 13.51 22.46 82.35 20.87
CCMSRNet [27] 22.74 88.74 13.18 22.22 89.07 10.76 18.96 77.67 23.85
DDFormer [29] 10.94 25.74 58.26 10.87 41.27 72.19 10.64 27.81 69.46
Semi-UIR [28] 23.63 81.81 23.08 19.83 73.95 30.54 19.12 74.94 25.68
UWFormer [30] 19.65 85.26 17.21 21.45 90.27 10.88 19.62 80.36 21.05
PATS-UIENet (Ours) 23.59 90.16 10.42 22.78 90.54 8.85 25.87 84.34 12.04
V-A2 Non-reference Quantitative Evaluation
We used two non-reference metrics, i.e., UIF [59] and MUSIQ [60], to assess the performance of our method and sixteen baselines on the Test-U80, Test-S110, Test-Scenes, Test-C60, Test-UCCS and Test-R53 testing sets. The results are shown in Table II. It can be seen that our method achieved the best UIF score on three out of the six testing sets, including Test-U80, Test-S110 and Test-C60, and ranked the second on the Test-R53 testing set. Regarding the MUSIQ metric, the proposed method still demonstrated the competitive performance, especially on the challenging Test-S110, Test-Scenes and Test-R53 testing sets, even though it did not always produce the highest score.
TABLE II: The non-reference quantitative evaluation of the proposed PATS-UIENet and sixteen baselines on six real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts, respectively.
Method Test-U80 Test-S110 Test-Scenes Test-C60 Test-UCCS Test-R53
[…]
UDCP [4] 36.40 44.99 32.01 51.38 61.25 36.99 25.41 33.03 34.72 29.72 3.24 36.85
DHL [10] 41.32 47.43 14.80 57.22 0.55 36.57 25.04 37.27 3.62 31.07 2.53 42.64
HE [9] 44.04 46.68 45.01 56.14 52.81 35.79 41.50 36.08 38.08 31.24 1.67 43.06
HLRP [15] 0.59 49.73 0.91 56.87 0.71 40.21 3.57 34.65 0.32 34.44 1.26 43.53
MMLE [14] 30.99 52.50 35.05 60.18 42.05 44.67 24.95 40.18 28.08 35.69 24.67 53.53
WaterNet [22] 33.25 31.84 28.50 40.47 72.94 38.47 28.65 34.75 33.38 24.88 9.71 32.93
UColor [23] 44.15 45.10 42.83 54.80 70.20 35.12 26.46 36.01 49.22 35.12 2.31 49.63
FUnIE-GAN [20] 29.12 47.47 22.77 52.27 37.69 45.38 23.11 43.49 22.67 34.64 10.93 38.92
USUIR [24] 5.17 44.44 37.02 56.83 48.82 37.91 22.35 34.15 20.54 28.83 17.98 39.44
PUIE-Net (MP) [25] 9.18 49.05 60.98 60.44 27.69 45.46 3.39 38.69 51.76 29.18 46.16 45.15
U-Transformer [21] 5.16 39.34 24.94 44.96 38.68 30.44 29.31 37.66 29.69 27.05 19.24 30.06
URanker [26] 54.87 44.63 58.31 57.15 72.54 39.20 44.70 38.37 45.97 29.60 39.19 41.92
CCMSRNet [27] 54.61 49.85 0.23 60.59 63.59 39.64 41.85 40.24 51.34 33.72 34.80 48.23
DDFormer [29] -5.36 36.99 1.92 40.68 -3.16 23.87 11.14 36.21 -2.14 20.19 5.86 24.87
Semi-UIR [28] 50.06 45.40 31.94 51.22 44.01 41.47 46.82 43.25 32.28 30.34 23.33 35.21
UWFormer [30] 46.48 50.95 55.69 61.53 67.40 36.80 36.76 41.40 53.24 34.30 39.98 46.26
PATS-UIENet (Ours) 58.83 49.91 61.20 60.66 68.07 46.70 46.99 40.49 51.18 31.63 42.51 46.63
V-A3 Qualitative Analysis
The enhanced images generated by the 16 baselines and our method on six testing sets are shown in Figs. 4 - 9, respectively. It can be seen that some methods with the higher UIF [59] or MUSIQ [60] score did not produce visually pleasing results. For example, MMLE [14] produced the highest MUSIQ score on Test-U80 and Test-R53, but the resulting images suffered from pale colors and over-enhancement artifacts (see Figs. 4 and 9). In contrast, our PATS-UIENet improved different degraded images and produced images with the natural and vivid color. Although the UIF [59] or MUSIQ [60] scores that our method produced were slightly lower than those obtained using some baseline methods [14, 22, 20] on certain testing sets, the enhanced images still manifest visually satisfactory quality.
Figure 4: The results produced by 16 baselines and our method in terms of a degraded image in the Test-U80 testing set.
Figure 5: The results produced by 16 baselines and our method in terms of a degraded image in the Test-S110 testing set.
Figure 6: The results produced by 16 baselines and our method in terms of a degraded image in the Test-Scenes testing set.
Figure 7: The results produced by 16 baselines and our method in terms of a degraded image in the Test-C60 testing set.
Figure 8: The results produced by 16 baselines and our method in terms of a degraded image in the Test-UCCS testing set.
Figure 9: The results produced by 16 baselines and our method in terms of a degraded image in the Test-R53 testing set.
V-A4 Performance Analysis
We compared the proposed PATS-UIENet with 11 deep baseline methods in terms of the number of parameters, Floating Point Operations Per Second (FLOPs) and inference speed (ms). Both the number of parameters and the FLOPs were calculated using the ptflops [^2] https://pypi.org/project/ptflops/ tool. As shown in Table III, the values of number of parameters, FLOPs and inference speed vary significantly. Specifically, UColor [23] exhibits the largest number of parameters (148.04M) and the largest FLOPs value (2804.36G), while UWFormer [30] incurs the slowest inference speed (491.00 ms). On the other hand, USUIR [24] has the lightest design with only 0.23M parameters and has a FLOPs value of 29.62G and owns the fastest inference speed (2.13ms). In contrast, our PATS-UIENet achieved a proper balance between the number of parameters or FLOPs and the inference speed.
TABLE III: Comparison of between 11 deep UIE methods and our approach in terms of number of parameters, FLOPs and inference speed.
Method #Params (M) FLOPs (G) Speed (ms)
WaterNet [22] 1.09 142.84 6.40
UColor [23] 148.04 2804.36 15.51
FUnIE-GAN [20] 7.02 20.48 73.18
USUIR [24] 0.23 29.62 2.13
PUIE-Net (MP) [25] 1.40 70.54 101.89
U-Transformer [21] 22.8 5.96 91.30
URanker [26] 3.15 20.90 45.50
CCMSRNet [27] 21.69 87.18 152.51
DDFormer [29] 7.63 35.54 50.04
Semi-UIR [28] 1.65 72.88 84.93
UWFormer [30] 29.84 30.18 491.00
PATS-UIENet (Ours) 45.71 179.70 86.17
V-B Transmission Estimation and Color Restoration
The quantitative evaluation of our methods and four prior-based baselines for transmission estimation and color restoration is reported in Table IV. It can be seen that our method achieved the superior, or at least the comparable, performance to that of the four prior-based baselines on four different subsets of Test-R53. Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets, no matter which metric was considered. It should be noted that none of these methods produced a PCC value higher than 0.12 on the Satil subset. The inferior results may be attributed to the special scenes contained in this subset [39]. As shown in Fig. 10, our method was able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
Figure 10: Each group shows a degraded image and the red channels of five transmission maps (top) and five enhanced images (bottom) obtained using four prior-based baselines and our method.
TABLE IV: The quantitative evaluation of four prior-based baselines and our method on the four subsets of Test-R53, where PCC and $\bar{\psi}$ are used for transmission estimation and color restoration, respectively.
Method Michmoret Katzaa Nachsholim Satil
[…]
DCP [3] -0.16 34.63 0.00 35.47 -0.26 34.75 0.03 36.21
UDCP [4] -0.54 36.97 -0.15 40.43 0.06 38.80 0.12 51.45
GDCP [5] 0.29 33.76 0.24 34.47 -0.04 34.28 -0.03 35.31
DHL [10] -0.07 32.10 0.10 35.50 0.36 35.92 0.11 34.82
Ours 0.62 12.83 0.32 9.47 -0.28 15.31 0.04 13.14
V-C Ablation Study
To investigate the effect of different components of the PATS-UIENet, we conducted a series of ablation experiments. For simplicity, we only utilized Test-U80 and the Michmorest subset in Test-R53.
V-C1 Effect of the Semi-supervised Learning Framework
To validate the effect of the proposed semi-supervised learning framework, we compared it with the unsupervised learning, supervised learning and bi-directional supervised learning frameworks. As reported in Table V, the proposed semi-supervised framework always outperformed the other frameworks in terms of different metrics across different testing sets. It has been suggested that our network can be better trained using the semi-supervised learning framework with both the labeled and unlabeled real-world images, which improves the enhancement performance.
TABLE V: Comparison of different learning frameworks on Test-U80 and the Michmorest subset in Test-R53.
Learning Framework Test-U80 Michmoret
[…]
Unsupervised 16.85 78.44 29.91 0.50 5.46
Supervised 22.19 88.23 13.07 0.57 17.01
Bi-supervised 23.00 89.49 11.93 0.61 15.02
Semi-supervised (Ours) 23.59 90.16 10.42 0.62 12.83
V-C2 Effect of the RCT and RCM
To examine the effect of the RCT and RCM on the performance of our PATS-UIENet, we removed the RCT, the RCM or both of them. In this case, we derived three variants of the PATS-UIENet: one with both the RCT and RCM removed, one with only the RCT removed and one with only the RCM removed, denoted as “Simplified”, “w/o RCT” and “w/o RCM”, respectively. As shown in Table VI, the complete PATS-UIENet achieved the best performance in terms of all the metrics. Removal of the RCT led to a slight increase in the PSNR and SSIM metrics compared to the Simplified version, but worsened the LPIPS and $\bar{\psi}$ scores. On the other hand, removal of the RCM reduced the performance with regard to all metrics across the two data sets, confirming its usefulness in facilitating information exchange between the B-Stream and A-Stream. The results demonstrate that both the RCT and RCM contribute meaningfully to the performance of our method.
TABLE VI: Impact of the RCT and RCM on the performance of our method on Test-U80 and the Michmorest subset in Test-R53.
Variant Test-U80 Michmoret
[…]
Simplified 22.96 89.64 11.25 0.59 13.59
w/o RCM 22.10 88.83 12.20 0.60 13.50
w/o RCT 23.14 90.03 12.20 0.60 13.97
Ours 23.59 90.16 10.42 0.62 12.83
V-C3 Effect of the Degradation Control Factor
To evaluate the impact of the degradation control factor $\alpha$ on the performance of the PATS-UIENet, we conducted an ablation experiment on testing three different values of $\alpha$. As shown in Table VII, the default setting of $\alpha$ = 0.1 achieved the best performance with regard to both the full-reference and the non-reference metrics across the two data sets.
TABLE VII: The effect of the degradation control factor, $\alpha$, on the performance of our method.
$\alpha$ Test-U80 Michmoret
[…]
0.001 23.21 89.77 11.02 0.59 13.92
1 23.06 89.72 11.08 0.61 14.11
0.1 (Ours) 23.59 90.16 10.42 0.62 12.83
V-C4 Effect of Loss Hyperparameters
To evaluate the impact of the hyperparameters used for the loss functions, including $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{unsup}$, on the performance of the proposed PATS-UIENet, we conducted a comprehensive ablation study by changing their values. Regarding the supervised learning loss function (see Eq. (7)), the results produced by our method with different combinations of the $\lambda_{1}$ and $\lambda_{2}$ values are shown in Table VIII. It can be seen that the combination that we used produced the best result in terms of each metric. For the unsupervised learning loss function (see Eq. (15)), the results produced by our method with different values of $\lambda_{3}$ are reported in Table IX. Again, the value that we chose led to the best result no matter what metric was considered. In terms of the semi-supervised learning loss function (see Eq. (16)), we present the results obtained using our method with different $\lambda_{unsup}$ values in Table X. As can be observed, the value that we utilized produced the best result with regard to each metric. The above findings highlight the effectiveness of the values of different loss hyperparameters that we chose.
TABLE VIII: The effect of different combinations of hyperparameters used for the loss function of the supervised learning scheme, i.e., $\lambda_{1}$ and $\lambda_{2}$, on the performance of our method.
$\lambda_{1}$ $\lambda_{2}$ Test-U80 Michmoret
[…]
1 0.0005 22.04 88.01 12.94 0.36 16.92
0.0001 0.0005 22.84 89.65 10.85 0.61 13.57
0.1 0.1 22.95 89.67 11.52 0.61 15.69
0.1 0 22.91 80.57 11.54 0.59 16.92
0.1 0.0005 (Ours) 23.59 90.16 10.42 0.62 12.83
TABLE IX: The effect of the weighting factor used for the loss function of the unsupervised learning scheme, i.e., $\lambda_{3}$, on the performance of our method.
$\lambda_{3}$ Test-U80 Michmoret
[…]
0.1 22.95 89.80 11.01 0.61 15.14
10 22.36 88.19 13.09 0.60 14.72
1 (Ours) 23.59 90.16 10.42 0.62 12.83
TABLE X: The effect of the weighting factor used for the loss function of the semi-supervised learning scheme, i.e., $\lambda_{unsup}$, on the performance of our method.
$\lambda_{unsup}$ Test-U80 Michmoret
[…]
1 22.32 88.75 12.16 0.61 14.08
0.01 23.31 89.96 10.92 0.61 12.84
0.1 (Ours) 23.59 90.16 10.42 0.62 12.83
V-C5 Effect of Different Stream Architectures
To further examine the effect of different stream architectures on the PATS-UIENet, we constructed its three variants by building the D-Stream, B-Stream and A-Stream using convolution and/or Transformer networks. In addition, we obtained a fourth variant by removing the D-Stream. In essence, this variant is equal to the network built on top of the original IFM [1]. The four variants were trained and tested using the same setup as that used for our PATS-UIENet. The results produced by these variants and our method are shown in Table XI. It can be seen that our method produced the better, or at least comparable, results in contrast to the four variants. This finding supports the design philosophy of the PATS-UIENet. That is to say, CNNs are suitable for capturing local characteristics while Transformer is good at encoding global characteristics. Moreover, our network which was built on top of the revised IFM [2] normally outperformed the variant constructed based on the original IFM [1], confirming the choice of the revised IFM [2].
TABLE XI: Comparison between our method and its four variants obtained by building the D-Stream, B-Stream and A-Stream using different networks.
Method Streams Test-U80 Michmoret
[…]
$\text{Variant}_{1}$ CNN CNN CNN 22.31 88.69 12.48 0.58 12.44
$\text{Variant}_{2}$ Trans Trans CNN 20.71 75.36 25.31 0.73 17.28
$\text{Variant}_{3}$ Trans Trans Trans 20.57 75.00 25.73 0.71 14.23
$\text{Variant}_{4}$ N/A CNN Trans 22.51 88.85 11.47 0.71 15.95
Ours CNN CNN Trans 23.59 90.16 10.42 0.62 12.83
VI Conclusion
In this paper, we introduced a novel Physics-Aware Triple-Stream Underwater Image Enhancement Network, namely, PATS-UIENet, which explicitly estimates the degradation parameters of the revised IFM, for the UIE task. To overcome the challenge of insufficient data, we also adopted an IFM-inspired semi-supervised learning framework, comprising a bi-directional supervised scheme and an unsupervised scheme. Due to the complementary action of the two schemes, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model is stronger, compared to supervised and unsupervised methods. To our knowledge, this study made the first effort to jointly exploit the physics-aware deep network and the IFM-inspired semi-supervised learning technique for the UIE task. The proposed method performed better than, or at least comparably to, sixteen baselines on six underwater testing sets in different evaluations. The promising results should be due to the fact that our method is able to not only estimate degradation parameters but also learn the characteristics of diverse underwater scenes.

### v6

@@
-Semi-supervised Underwater Image Enhancement Using A Physics-Aware Triple-Stream Network
+Physics-Aware Semi-Supervised Underwater Image Enhancement
 Abstract
@@
 However, the inflexible assumption of the former often impairs their effectiveness in handling diverse underwater scenes, while the generalization of the latter to unseen images is usually weakened by insufficient data.
-In this study, we leverage both the physics-based Image Formation Model (IFM) and deep learning techniques for Underwater Image Enhancement (UIE).
-To this end, we propose a novel Physics-Aware Triple-Stream Underwater Image Enhancement Network, i.e., PATS-UIENet, which comprises a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Light Estimation Stream (A-Stream).
-This network fulfills the UIE task by explicitly estimating the degradation parameters of a revised IFM.
+In this study, we leverage both the physics-based underwater Image Formation Model (IFM) and deep learning techniques for Underwater Image Enhancement (UIE).
+To this end, we propose a novel Physics-Aware Dual-Stream Underwater Image Enhancement Network, i.e., PA-UIENet, which comprises a Transmission Estimation Steam (T-Stream) and an Ambient Light Estimation Stream (A-Stream).
+This network fulfills the UIE task by explicitly estimating the degradation parameters of the IFM.
 We also adopt an IFM-inspired semi-supervised learning framework, which exploits both the labeled and unlabeled images, to address the issue of insufficient data.
-To our knowledge, such a physics-aware deep network and the IFM-inspired semi-supervised learning framework have not been used for the UIE task before.
-Our method performs better than, or at least comparably to, sixteen baselines across six testing sets in the degradation estimation and UIE tasks.
-These promising results should be due to the fact that the proposed method can not only model the degradation but also learn the characteristics of diverse underwater scenes.
-[^1] The models and code will be published on the acceptance of the paper.
-Index Terms:
-Underwater Image Enhancement (UIE), Underwater Image Processing, Image Formation Model (IFM), Deep learning, Semi-supervised Learning.
-I Introduction
+Our method performs better than, or at least comparably to, eight baselines across five testing sets in the degradation estimation and UIE tasks.
+This should be due to the fact that it not only can model the degradation but also can learn the characteristics of diverse underwater scenes.
+1 Introduction
 The images captured in the underwater environment play important roles in ocean exploration.
 However, these images normally suffer from different degradation, due to the wavelength-dependent light absorption and light scattering caused by the underwater transmission medium.
-According to the Image Formation Model (IFM) [1], the process of image degradation can be formulated as follows:
+According to [1], the underwater Image Formation Model (IFM) can be expressed as:
 $I^{c}(x)=J^{c}(x)t^{c}(x)+(1-t^{c}(x))A^{c},$ (1)
 where $I^{c}(x)$ ($c\in\{R,G,B\}$) is a degraded image, $J^{c}(x)$ is the underlying clean image, $t^{c}(x)=exp(-\beta^{c}d(x))$ denotes the transmission maps, $\beta^{c}$ is the attenuation factor for each channel, $d(x)$ is the scene distance map and $A^{c}$ is the ambient light.
-(See Fig.
-1 for examples).
+(See Fig.1 for examples).
 In this context, $J^{c}(x)t^{c}(x)$ describes the distance-dependent color distortion, while $(1-t^{c}(x))A^{c}$ represents the scattering of the ambient light which decreases the contrast and lessens the visibility of the image.
-Despite the IFM had been widely used, its flaws in accurately modeling the degradation process was highlighted [2].
-Figure 1: Examples of a degraded underwater image, the corresponding ambient light image $A^{c}$ and three transmission maps $t^{c}(x)$ ($c\in\{R,G,B\}$) estimated, and the enhanced image produced using the estimated parameters of the IFM [1].
-Considering that the degradation may interfere with downstream tasks, the design of effective UIE methods is critical yet challenging.
-Existing methods can be divided into two categories, i.e., traditional prior-based methods and deep learning-based methods.
-Traditional prior-based methods [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] typically rely on a certain form of prior information.
-Some prior-based methods [3, 4, 5] aim at estimating the parameters of the IFM.
+Figure 1: Examples of a degraded underwater image, the corresponding ambient light image $A^{c}$ and three transmission maps $t^{c}(x)$ ($c\in\{R,G,B\}$) estimated using our method, and the enhanced image produced using the estimated parameters on top of the IFM.
+Considering the degradation may interfere with downstream tasks, Underwater Image Enhancement (UIE) is key to underwater image processing and machine vision.
+Due to the complicated degradation process, this task is challenging.
+Many traditional methods [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] have been proposed, to address the challenge.
+These methods can be treated as the prior-based methods because they typically rely on a certain form of prior information.
+Some prior-based methods [2, 3, 4] aimed to estimate the parameters of the IFM.
 However, they normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene violates the prior assumption [4].
-On the other hand, the other prior-based methods [11, 12, 13, 14, 15] directly enhance the quality of underwater images without explicitly modeling the degradation process.
-Due to the lack of the knowledge of underwater scenes and the limited adaptability and robustness, they cannot effectively enhance the images captured in the complex scenes and often introduce artifacts.
-In contrast, deep learning-based UIE approaches [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] aim to directly learn from different types of degraded images.
-Therefore, the problem with the limited static priors can be alleviated.
+On the other hand, the other prior-based methods [9, 10, 11] directly enhance the quality of underwater images without explicitly modeling the degradation process.
+Due to the lack of the knowledge of underwater scenes, they cannot effectively enhance the images captured in the complex scenes and often introduce artifacts.
+In contrast, deep learning-based UIE approaches [14, 15, 16, 17] aim to directly learn from different types of degraded images.
+Therefore, the problem with the limited static prior can be alleviated.
 Ideally, a large number of real-world labeled training images should be utilized.
@@
 In this case, the deep UIE model trained may overfit the training data.
-The synthetic images generated using the IFM [32, 18] or a Generative Adversarial Network (GAN) [33, 34, 35, 20] were used to address the problem with the lack of labeled data.
+The synthetic images generated using the IFM [19, 14] or a Generative Adversarial Network (GAN) [20, 21, 22] were employed in order to address the problem with the lack of labeled data.
 However, the utilization of these images to simulate real-world underwater degradation processes is struggling.
@@
 But they are difficult to train and may produce unstable results.
-Although unsupervised learning methods [36, 37] alleviate the reliance on the paired data by directly learning degradation parameters from unlabeled degraded images, they often suffer from slow inference speed and are prone to color distortion.
-Recently, semi-supervised learning methods [38, 28, 29, 30], which combine the strengths of both the supervised and unsupervised approaches, have achieved promising results by leveraging the labeled data along with the unlabeled data during the training process.
-Despite the progress has been made by the existing methods, two challenges remain.
-First, existing deep learning-based methods lack effective physics-aware modeling, which limits their interpretability and effectiveness in modeling the degradation process.
-Second, a supervised or unsupervised approach normally struggles with the insufficient or noisy data due to the complexity of real underwater degradation, leading to overfitting or weak generalization.
-To address these challenges, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network, i.e., PATS-UIENet, which explicitly incorporates the revised IFM [2] into the deep neural network that we deliberately design.
-Unlike existing physics-based methods [3, 4, 5], our network explicitly estimates the parameters of a physical model using three streams, including a Direct Signal Transmission Estimation Stream (D-Stream), a Backscatter Signal Transmission Estimation Stream (B-Stream) and an Ambient Estimation Stream (A-Stream).
-Due to the explicit modeling of the physical degradation process, our network is superior to existing UIE methods by exploiting the strengths of both the theoretical interpretability of physics-based approaches and the powerful feature representation ability of deep neural networks.
-We further introduce an IFM-inspired semi-supervised learning framework, including a bi-directional supervised scheme and an unsupervised scheme, for the purpose of overcoming the limitation of insufficient data.
-The supervised scheme effectively utilizes the limited labeled data by providing explicit degradation guidance, while the unsupervised scheme exploits the abundant unlabeled data to improve generalization.
-Therefore, our framework can be better trained than the existing supervised approaches [18, 19, 22, 23, 24, 25, 26, 27], unsupervised methods [36, 37] and semi-supervised approaches [38, 28, 29, 30] using the same amount of data.
-As a result, the generalization of the model trained using our framework is stronger than that trained using existing supervised, unsupervised and semi-supervised approaches.
-To our knowledge, this study makes the first effort to jointly apply the physics-aware deep network and the IFM-inspired semi-supervised learning technique to the UIE task.
+Recently, promising results were derived using zero-shot learning approaches [23, 24] by directly learning degradation parameters from unlabeled degraded images.
+To effectively model the degradation process, we propose a Physics-Aware Dual-Stream Underwater Image Enhancement Network, i.e., PA-UIENet, by injecting the physical principle into the deep neural network.
+This network performs the UIE task by explicitly estimating the degradation parameters of the IFM using a Transmission Estimation Steam (T-Stream) and an Ambient Light Estimation Stream (A-Stream).
+We further adopt an IFM-inspired semi-supervised learning framework, including a bi-directional supervised scheme and an unsupervised scheme, to overcome the limitation of insufficient data.
+Compared with the unsupervised methods [23, 24], the framework can be more easily trained due to the facilitation from the supervised scheme.
+Since the unsupervised scheme uses additional unlabeled images while the bi-directional supervised scheme exploits labeled images more effectively, it can be better trained than the supervised approaches [14, 15, 16, 17].
+As a result, the generalization of the model trained using our framework is stronger than that trained using a supervised or unsupervised method.
+To our knowledge, this study makes the first attempt to jointly apply the physics-aware deep network and the IFM-inspired semi-supervised learning technique to the UIE task.
 Our main contributions can be summarized as threefold.
-  • We introduce a Physics-Aware Triple-Stream Underwater Image Enhancement Network, referred to as PATS-UIENet, which can explicitly estimate the degradation parameters of the revised IFM.
-This network differentiates itself from existing IFM-based methods by integrating physical parameter estimation into a deliberately designed deep neural network.
-Such an IFM-motivated network has not been explored for the UIE task before.
-  • We propose an IFM-inspired semi-supervised learning framework, which addresses the issues of data insufficiency and training instability by exploiting the merits of both the supervised and unsupervised learning methods.
-  • We conduct a series of comparative experiments on six underwater testing sets along with sixteen baselines.
-The results not only validate the effectiveness of our method but also provide benchmarks for future research.
-The remainder of this paper is organized as follows.
-The related literature is reviewed in Section II.
-We introduce the proposed method in Section III.
-The experimental settings and results are reported in Sections IV and V respectively.
-Finally, our conclusion is drawn in Section VI.
-II Related Work
-II-A Revised Image Formation Model
-As pointed out by Akkaynak and Treibitz [2], the original IFM [1] ignored the dependencies of the backscatter coefficient on the ambient light and the optical property of water bodies.
-They assumed that the attenuation coefficients of the direct signal and the backscattered signal are the same.
-In this case, the original IFM [1] cannot adequately describe the degradation process of underwater images.
-It has been demonstrated that the attenuation factors of the direct signal and the backscatter signal are different while the attenuation factor of the backscatter signal is affected more severely by the ambient light [2].
-Akkaynak and Treibitz [2] further proposed a revised IFM, which can be expressed as:
-$I^{c}(x)=J^{c}(x)e^{-\beta_{c}^{D}(\mathbf{D})\cdot z}+\left(1-e^{-\beta_{c}^{B}(\mathbf{B})\cdot z}\right)A^{c},$ (2)
-where $\textbf{D}=\{z,\rho,E,S_{c},\beta\}$ and $\textbf{B}=\{E,S_{c},b,\beta\}$ are two sets of parameters which affect the attenuation factors of the direct signal $\beta_{c}^{D}$ and the backscatter signal $\beta_{c}^{B}$, respectively, $z$ represents the distance between the scene and the camera, $\rho$ is the reflectance spectrum of the object, $E$ is the ambient light at a certain distance, $S_{c}$ is the camera response function, and $b$ and $\beta$ are the beam scattering and attenuation coefficients, respectively.
-In contrast to the original IFM [1], the revised IFM [2] offers the more comprehensive and physically accurate representation of underwater image degradation.
-Therefore, we design the proposed method on top of the revised IFM.
-For more details, please refer to the original publication [2].
-II-B Prior-Based Methods
+(1) We propose a physics-aware dual-stream network for the UIE task by explicitly estimating the degradation parameters of the IFM.
+(2) An IFM-inspired semi-supervised learning framework is adopted, to overcome the challenge of insufficient data.
+(3) We conduct a series of comparative experiments on five real-world underwater testing sets along with eight baselines.
+These results provide the community with a set of benchmarks.
+2 Related Work
+2.1 Prior-Based Methods
 To recover clear underwater images, many methods [2, 3, 4, 5, 6, 7, 8] used the prior information to estimate the degradation parameters of the IFM.
@@
 Moreover, the Haze Line Prior (HL) [7] was utilized in some studies [8, 25].
-Despite these methods were designed on top of the IFM, they normally made a rigid assumption about the underwater environment, which restricted the application of them to a specific underwater scenario.
-On the other hand, some approaches used the more general priors [40, 11, 12, 13, 16, 14, 15] to enhance the quality of underwater images without taking the IFM into account, including Histogram Equalization (HE) [40, 11], Retinex-based methods [13, 16] and image fusion techniques [17, 41].
+Despite these methods were designed on top of the IFM, they normally made a rigid assumption about the underwater environment, which restricted the application of these to a specific underwater scenario.
+On the other hand, some approaches used the more general priors [26, 9, 10, 11, 12] to enhance the quality of underwater images without taking the IFM into account, including Histogram Equalization (HE) [26, 9], Retinex-based methods [11, 12] and image fusion techniques [13, 27].
 These approaches usually enhanced the contrast and produced more color-balanced results.
 However, they probably struggle with processing globally inhomogeneous degraded images and may introduce artifacts, such as halos and color casts, due to the lack of the knowledge of underwater scenes.
-II-C Learning-Based Methods
-Thanks to the powerful representation learning ability and large-scale training data [42], deep learning-based methods greatly boosted the development of computer vision in both high-level [43, 44, 45] and low-level vision tasks [46, 47, 48].
+2.2 Learning-Based Methods
+Thanks to the powerful representation learning ability and large-scale training data [28], deep learning based methods greatly boosted the development of computer vision in both high-level [29, 30, 31] and low-level vision tasks [32, 33, 34].
 However, the application of deep learning to underwater image enhancement is much less than other tasks.
@@
 For example, the WaterNet [16] used CNNs to learn fusion weights for the purpose of fusing the enhanced results produced by the White Balance [13], HE and Gamma Correction techniques.
-To perform the UIE task, the UColor [23] method combined the RGB, HSV and Lab color spaces with the guidance of the IFM parameters estimated using the GDCP [5].
-Since the training data sets were relatively small, these methods usually overfitted the training data and poorly generalized to unseen images.
-In addition, they could not explicitly utilize the information provided by the IFM.
+To perform the UIE task, the UColor [17] method combined the RGB, HSV and Lab color spaces with the guidance of the IFM parameter estimated using the GDCP [4].
+Since the training data sets were relatively small, these methods usually overfitted the training data and were poorly generalized to unseen images.
+In addition, they could not explicitly exploit the information provided by the IFM.
 In contrast to the lack of labeled underwater images, it is easier to collect unlabeled underwater images.
 Existing studies [15, 36] sourced real-world underwater images and categorized them according to the degree of degradation.
-Unsupervised learning methods [24, 36] were developed on top of these data by learning the mapping from severely degraded images to slightly degraded images.
-However, these approaches were difficult to train and tended to produce unstable results.
-Recently, semi-supervised learning methods [38, 28, 29, 30] have been developed, using both labeled and unlabeled underwater images.
-Although promising results were derived, these methods did not consider physical principles.
+Unsupervised learning approaches were developed on top of these data by learning the mapping from severely degraded images to slightly degraded images.
+However, these approaches were hard to train and tended to produce unstable results.
+Recently, a semi-supervised UIE method [37] was developed.
+Nevertheless, this method did not consider physical principles.
 The above-mentioned studies either are not robust to diverse underwater scenes, or lack the knowledge of underwater images.
-We are hence motivated to exploit both the IFM-inspired semi-supervised learning technique and the physics-aware deep neural network, to address these issues.
-III Methodology
-Considering the importance of physical principles to the UIE task, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network (PATS-UIENet).
-This network contains a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Light Estimation Stream (A-Stream), which are used to explicitly estimate the three parameters of the revised IFM [2], respectively.
+We are hence inspired to exploit both the IFM-inspired semi-supervised learning technique and the physics-aware deep neural network to address these issues.
+3 Methodology
+Considering the importance of physical principles to the UIE task, we propose a Physics-Aware Dual-Stream Underwater Image Enhancement Network, namely, PA-UIENet.
+This network contains a Transmission Estimation Steam (T-Stream) and an Ambient Light Estimation Stream (A-Stream), which are used to explicitly estimate the parameters of the IFM.
 To address the challenge of the lack of labeled real-world underwater images, we further adopt an IFM-inspired semi-supervised learning framework, which consists of a bi-directional supervised scheme and an unsupervised scheme.
-Compared to the supervised or unsupervised method, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model trained is thus stronger, due to the complementary action of the two schemes.
-III-A Physics-Inspired Design
-Motivated by the revised IFM [2], we propose a physics-aware underwater image enhancement network on top of this model.
-Since the network explicitly estimates the parameters of the revised IFM by learning from training images, it avoids the rigid constraints of priors while improving the interpretability of the model trained.
-To reduce the complexity of the network, we simplify Eq.
-(2) as follows:
-$I^{c}(x)=J^{c}(x)t_{D}^{c}(x,\textbf{D})+\left(1-t_{B}^{c}(x,\textbf{B})\right)A^{c}.$ (3)
-where $I^{c}(x)$ ($c\in\{R,G,B\}$) is the degraded image, $J^{c}(x)$ is the underlying clean image, $t_{D}^{c}(x,\textbf{D})$ and $t_{B}^{c}(x,\textbf{B})$ represent the direct and backscatter signal transmission maps, respectively, and $A^{c}$ denotes the ambient light.
-To estimate the three parameters, including $t_{D}^{c}(x,\textbf{D})$, $t_{B}^{c}(x,\textbf{B})$ and $A^{c}$, we design a D-Stream, a B-Stream and an A-Stream, respectively.
-We also introduce an IFM-inspired semi-supervised learning framework by leveraging the explicitly estimated parameters of the revised IFM [2].
-This framework consists of a bi-directional supervised learning scheme and an unsupervised learning scheme, which addresses the challenge of rare labeled real-world underwater images.
-Compared to the supervised or unsupervised methods, our PATS-UIENet can take advantage of both labeled and unlabeled real-world images for the more effective training operation.
-Furthermore, the complementary action between the two schemes enhances the generalization ability of the model trained.
-III-B PATS-UIENet
-As illustrated in Fig.
-2, the PATS-UIENet contains two encoder-decoder style streams with the same structure, i.e., D-stream and B-stream, and a Transformer-based A-stream.
-The three streams are used to estimate the direct signal transmission map, the backscatter signal transmission map and the ambient light, respectively.
-Since the red channel usually encounters more severe attenuation than the blue and green channels [1, 51], the utilization of this channel is key to color restoration.
-Therefore, a simple Red Channel Tuner (RCT) module is designed to adaptively emphasize the red channel before an image is fed into the transmission estimation steams.
-The output of the first encoder block $Enc\_1$ in the B-stream is processed by Patchify and the result is sent to the A-stream.
-A Residual Communication Module (RCM) connects the two blocks of the encoder of the B-stream and the A-stream at the same level, to fulfill the bi-directional feature exchange.
-The two sets of transmission maps estimated using the D-stream and B-stream, respectively, and the ambient light estimated using the A-stream are fed into the revised IFM [2].
+Compared to the supervised or unsupervised method, the PA-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model trained is thus stronger, due to the complementary action of the two schemes.
+Figure 2: The architecture of the proposed PA-UIENet, which comprises two individual streams: A-Stream and T-Stream, to estimate the degradation parameters of the IFM.
+3.1 PA-UIENet
+As illustrated in Fig.2, the PA-UIENet contains an encoder-decoder style T-stream and a Transformer-based A-stream, used to estimate transmission maps and the ambient light respectively.
+Since the red channel usually encounters the more severe attenuation than the blue and green channels [1, 38], the utilization of this channel is key to color restoration.
+Therefore, a simple Red Channel Tuner (RCT) module is designed to adaptively emphasize the red channel before an image is fed into the T-stream.
+The output of the first encoder block Enc_1Enc\_1 in the T-stream is processed by Patchify and the result is sent to the A-stream.
+A Residual Communication Module (RCM) connects the two blocks of the encoder of the T-stream and the A-stream at the same level, to fulfill the bi-directional feature exchange.
+Both the transmission maps and the ambient light, estimated using the T-stream and A-stream respectively, are fed into the IFM.
 As a result, the enhanced underwater image is produced.
-Figure 2: The architecture of the proposed PATS-UIENet, which comprises three individual streams, namely, D-Stream, B-Stream and A-Stream, to estimate the degradation parameters of the revised IFM [2].
-Red Channel Tuner (RCT).
+Red Chanel Tuner (RCT).
 Inspired by the channel attention mechanism [39], we propose a compact RCT module in order to adaptively emphasize the information contained in the red channel.
@@
 The weight is finally used to scale the red channel of the image.
-D-Stream.
-D-stream is designed to estimate the direct signal transmission map of an underwater image.
+T-Stream.
 According to Eq.
-(3), the transmission map varies at the locations of different pixels.
+(1), the transmission map varies at different pixel locations.
 In this case, the network should be able to generate fine-grained representations.
-Hence, we build the D-Stream using a CNN-based encoder-decoder network, due to its strong ability to learn local characteristics, which helps preserve edges and fine details.
-The encoder comprises five consecutive blocks, denoted as $Enc\_1$ to $Enc\_5$.
+Hence, we build the T-stream using a CNN-based encoder-decoder network.
+The encoder comprises five blocks, denoted as Enc_1Enc\_1 to Enc_5Enc\_5, and the Max Pooling operation is applied to each block.
 As a result, multi-scale feature maps are derived using the encoder.
-Given $Enc\_5$ is used as the bottleneck, the decoder contains four blocks, denoted as $Dec\_1$ to $Dec\_4$, symmetrically.
+Given Enc_5Enc\_5 is used as the bottleneck, the decoder contains four blocks, denoted as Dec_1Dec\_1 to Dec_4Dec\_4, symmetrically.
+Bilinear interpolation is used to upsample the feature maps within each block.
+Both the encoder and decoder blocks share the same structure, which consists of two ConvConv-BNBN-ReLUReLU operations.
 Skip connection is used to pass the feature maps produced by an encoder block to the corresponding decoder block at the same level, which is useful for restoring the fine-grained spatial structure.
-The use of multi-scale features and skip connections enhances the ability of the network to restore the spatial structure of the scene.
 The output of the last decoder block is fed into a convolutional layer.
-Three transmission maps are produced in terms of different color channels.
-In essence, D-Stream is specifically focused on modeling the directly transmitted light, which normally contains the spatial structure of the scene.
-B-Stream.
-B-stream is responsible for modeling the degradation caused by the backscattering in underwater environments.
-Although it shares the same CNN-based architecture as D-Stream for training stability, B-stream differs in functionality and processing strategy.
-Unlike direct signals, backscatter signals exhibit strong global interference and require a context-aware estimation.
-To this end, we pass the output of the first encoder block $Enc\_1$ through a Patchify processing and feed it into the A-Stream, leveraging the global modeling capability of Transformer to enhance the estimation accuracy.
-Moreover, B-Stream is connected with the A-Stream via Residual Communication Modules (RCMs), enabling bidirectional feature exchange.
-Compared to the D-Stream, B-Stream is focused on modeling the global degradation rather than recovering the local scene content.
+A set of transmission maps are produced in terms of different color channels.
 A-Stream.
-Ambient light reflects the overall luminance in the underwater scene and is generally spatially homogeneous.
-To effectively model this global characteristic, we employ a Transformer-based design instead of CNNs, because Transformer [53] is well-suited for capturing long-range dependencies and global characteristics through the self-attention mechanism.
-A-Stream consists of five Transformer blocks, denoted as $Trans\_1$ to $Trans\_5$, and includes a dedicated “Ambient” token to specifically learn the ambient light parameter.
-Additionally, A-Stream is coupled with the B-Stream through the RCM, allowing it to incorporate local features and mitigate the lack of spatial detail inherent in Transformer.
-A-Stream is distinguished from the other streams by its global feature learning process and Transformer-based architecture, which complements the local characteristic modeling of the CNN-based streams.
+The ambient light describes the luminance of the scene.
+Since this parameter does not encode the information of a specific object, it can be treated as homogeneous.
+We are motivated to design the A-stream using a set of Transformer [40] blocks for the sake of leveraging the self-attention mechanism to extract the global information.
+Five blocks, denoted as Trans_1Trans\_1 to Trans_5Trans\_5, are comprised of this stream.
+Each block contains a self-attention layer and a two-layer Multi-layer Perceptron (MLP) unit, in which two residual connections are used respectively.
+Since the input of the Transformer is a sequence of visual tokens , the feature maps of Enc_1Enc\_1 are processed using Patchify which includes the Flatten and Linear Projection operations.
+Considering the estimation of the ambient light does not require local characteristics and the quadratic complexity of self-attention, we downsample these maps by a factor of 16 to reduce the computational demand.
+The result is a set of tokens.
+A special token, namely, the Ambient token, is also adopted, to specifically learn the ambient light parameter.
+All the tokens are fed into Trans_1Trans\_1.
+The Ambient token produced by the last block Trans_5Trans\_5 is sent to a linear layer.
+Finally, the estimation of the ambient light is achieved.
 Residual Communication Module (RCM).
-According to the revised IFM [2], the degradation can be considered as the outcome caused by both the transmission and the ambient light.
-In this situation, the backscatter signal transmission estimation steam and the ambient light estimation stream will learn some common features.
+According to the IFM, the degradation can be considered as the outcome caused by both the transmission and the ambient light.
+In this situation, the T-stream and A-stream will learn some common features.
 Therefore, the feature exchange between both the streams is likely to boost the training of each stream.
@@
 We are inspired to propose the RCM for the sake of bridging the two streams.
-Specifically, the tokens produced by a block in the A-stream are folded into a set of feature maps while the feature maps generated by a block of the encoder of the B-stream are downsampled to the shape of these maps.
+Specifically, the tokens produced by a block in the A-stream are folded into a set of feature maps while the feature maps generated by a block of the encoder of the T-stream are downsampled to the shape of these maps.
 Then both sets of maps are concatenated along the channel dimension and are fed into a convolutional layer.
@@
 Each set is added with the original feature maps of the related stream.
-III-C IFM-Inspired Semi-supervised Learning Framework
-To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig.
-3), inspired by the revised IFM [2] theory, embedding physical constraints directly into the learning process to effectively leverage unlabeled data.
+3.2 IFM-Inspired Semi-Supervised Learning Framework
+To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig.3), inspired by the IFM theory.
 This framework includes a bi-directional supervised scheme and an unsupervised scheme.
-The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated parameters of the revised IFM to perform UIE, and a backward-degradation module, which uses the estimated parameters to degrade the reference images to their real-world degraded counterparts.
+The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated IFM parameters to perform UIE, and a backward-degradation module, which uses estimated parameters to degrade the reference images to their real-world degraded counterparts.
 As a result, the bi-directional supervised scheme can learn a more accurate estimation of the parameters using limited real-world data than that learned using a supervised method with a large number of synthetic data.
@@
 On the other hand, the unsupervised scheme exploits the unlabeled real-world underwater images that we collect.
-A second set of images with different degrees of degradation are generated from these images based on the revised IFM [2].
-Two sets of parameters can be estimated using both sets of images, respectively.
+A second set of images with different degrees of degradation are generated from these images based on the IFM.
+Two sets of parameters can be estimated using both sets of images respectively.
 Each set is utilized as the reference for the other set because their IFMs are related.
-Figure 3: The proposed semi-supervised learning framework, which comprises a bi-directional supervised learning scheme and an unsupervised learning scheme, used for training our PATS-UIENet.
+Figure 3: The proposed semi-supervised learning framework, which comprises a bi-directional supervised learning scheme and an unsupervised learning scheme, used for training our PA-UIENet.
 Bi-directional Supervised Learning.
@@
 Then we perform the forward-enhancement learning module by minimizing the following loss function:
-$\mathcal{L}_{fwd}=||(J^{c}(x)-\hat{J}^{c}(x))||_{2}^{2}.$ (4)
-Due to the ill-posed nature of the revised IFM, however, only $\mathcal{L}_{fwd}$ may be insufficient for ensuring the validity of the estimated parameters.
+[…]
+Due to the ill-posed nature of the IFM, however, only $\mathcal{L}_{fwd}$ may be insufficient for ensuring the validity of the estimated parameters.
 Therefore, we propose a backward-degradation learning module which degrades clear reference images using the estimated parameters to their degraded counterparts $\hat{I}^{c}(x)$.
 The supervision over the backward-degradation is achieved by minimizing the loss function:
-$\mathcal{L}_{bwd}=||(I^{c}(x)-\hat{I}^{c}(x))||_{2}^{2}.$ (5)
+[…]
 Inspired by the Retinex [11] theory, we further apply Gaussian blur to input images in order to derive the hint of the global ambient light, which is able to boost the training of the A-stream.
@@
 Unsupervised Learning.
-To leverage unlabeled underwater images, we adopt an unsupervised learning scheme based on the physical modeling capability of the revised IFM [2] rather than using generative networks [34, 19], which are difficult to train and often produce unstable results.
-This scheme adaptively constructs pairs of images with different levels of degradation from the unlabeled data.
-Given an unlabeled real-world degraded image $I^{c}_{1}(x)$, the revised IFM can be expressed as:
-$I^{c}_{1}(x)=J^{c}_{1}(x)t_{D1}^{c}+(1-t_{B1}^{c}(x))A^{c}_{1}.$ (8)
-A new degraded image $I^{c}_{2}(x)$ can be purposely derived according to:
-$I^{c}_{2}(x)=\alpha I^{c}_{1}(x)+(1-\alpha)(1-t_{B1}^{c}(x))A^{c}_{1},$ (9)
+To leverage unlabeled underwater images, generative networks [21, 15] were used.
+However, they were difficult to train and produced unstable results.
+Since unsupervised learning methods [23, 24] are able to directly learn from degraded images without requiring reference images, we are motivated to adopt an unsupervised learning scheme on top of the IFM.
+Given an unlabeled real-world degraded image $I^{c}_{1}(x)$, the IFM can be expressed as:
+$I^{c}_{1}(x)=J^{c}_{1}(x)t^{c}_{1}(x)+(1-t^{c}_{1}(x))A^{c}_{1}.$ (6)
+As performed in [23], a new degraded image $I^{c}_{2}(x)$ can be purposely derived according to:
+$I^{c}_{2}(x)=\alpha I^{c}_{1}(x)+(1-\alpha)A^{c}_{1},$ (7)
 where $\alpha\in(0,1)$ is a controlled factor which decides the degradation extent. Substituting $I^{c}_{1}(x)$ in Eq. (6) into Eq. (7), $I^{c}_{2}(x)$ can be then expressed as:
-$I^{c}_{2}(x)=J^{c}_{1}(x)(\alpha t_{D1}^{c}(x))+(1-t_{B1}^{c}(x))A^{c}_{1}.$ (10)
+$I^{c}_{2}(x)=J^{c}_{1}(x)(\alpha t^{c}_{1}(x))+(1-\alpha t^{c}_{1}(x))A^{c}_{1}.$ (8)
 Eq.
-(10) also satisfies the revised IFM but presents a more severe degradation than the $I^{c}_{1}(x)$, generated from the same underlying clear image $J^{c}_{1}(x)$, because $\alpha t_{D1}^{c}(x)<t_{D1}^{c}(x)$ always holds true.
-Therefore, the parameters $\hat{t}_{D2}^{c}(x)$ and $\hat{t}_{B2}^{c}(x)$ estimated using the PATS-UIENet from image $I^{c}_{2}(x)$ will be close to $\alpha\hat{t}_{D1}^{c}(x)$ and $\hat{t}_{B1}^{c}(x)$, respectively, where $\hat{t}_{D1}^{c}(x)$ and $\hat{t}_{B1}^{c}(x)$ are the parameters estimated from $I^{c}_{1}(x)$.
+(8) also satisfies the underwater IFM but presents the more severe degradation than the $I^{c}_{1}(x)$, generated from the same underlying clear image $J^{c}_{1}(x)$, because $\alpha t^{c}_{1}(x)<t^{c}_{1}(x)$ always holds true.
+Therefore, the parameters $\hat{t}^{c}_{2}(x)$ and $\hat{A}^{c}_{2}$ estimated using the PA-UIENet from image $I^{c}_{2}(x)$ will be close to $\alpha\hat{t}^{c}_{1}(x)$ and $\hat{A}^{c}_{1}$ respectively, where $\hat{t}^{c}_{1}(x)$ and $\hat{A}^{c}_{1}$ are the parameters estimated from $I^{c}_{1}(x)$ [23].
 Correspondingly, two loss functions are defined as:
-$\mathcal{L}_{D}=||\hat{t}_{D2}^{c}(x)-\alpha\hat{t}_{D1}^{c}(x)||_{2}^{2},$ (11)
-$\mathcal{L}_{B}=||\hat{t}_{B1}^{c}(x)-\hat{t}_{B2}^{c}(x)||_{2}^{2}.$ (12)
-Similar to Eq. (6), the loss function for the ambient light $\hat{A}^{c}$ estimated by the A-stream is defined as:
-$\mathcal{L}_{A\text{-}unsup}=||(I^{c}(x)*G-\hat{A}^{c}||_{2}^{2}.$ (13)
+[…]
+[…]
 Due to the lack of reference images, we use a non-reference loss function [24] based on the Gray-world prior [42] to constrain the enhancement of image $\hat{J}^{c}_{1}(x)$, which is expressed as:
-$\mathcal{L}_{gw}=\displaystyle\sum_{c\in\{R,G,B\}}||E(J^{c}_{1}(x))-0.5||_{2}^{2},$ (14)
+[…]
 where $E(J^{c}_{1}(x))$ denotes the average of a specific channel $c$. Finally, we define the loss function of the unsupervised learning scheme as:
-$\mathcal{L}_{unsup}=\mathcal{L}_{D}+\mathcal{L}_{B}+\mathcal{L}_{A\text{-}unsup}+\lambda_{3}\mathcal{L}_{gw},$ (15)
+$\mathcal{L}_{unsup}=\mathcal{L}_{T}+\mathcal{L}_{A\text{-}unsup}+\lambda_{3}\mathcal{L}_{gw},$ (12)
 where $\lambda_{3}$ is a weighting factor.
@@
 where $\lambda_{unsup}$ is used to balance the two schemes.
-IV Experimental Settings
+4 Experimental Settings
 In this section, we will briefly introduce the baselines, data sets, evaluation metrics and implementation details utilized in our experiments.
-IV-A Baselines
-We compared the proposed method with two prior-based approaches which were developed on top of the IFM [1], including UDCP [4] and DHL [10], three prior-based methods which did not consider the IFM, including Histogram Equalization (HE) [9], HLRP [15] and MMLE [14].
-We also compared our method with seven supervised learning approaches, including WaterNet [22], UColor [23], FUnIE-GAN [20], PUIE-Net (MP) [25], U-Transformer [21], URanker [26] and CCMSRNet [27], an unsupervised learning method, i.e., USUIR [24], and three semi-supervised learning methods, including DDFormer [29], Semi-UIR [28] and UWFormer [30].
-IV-B Data Sets
-We conducted a series of experiments using five data sets, including four publicly available real-world underwater data sets, i.e., UIEB [22], SUIM-E [31], RUIE [56] and SQUID [10], and a synthetic data set, namely, EUVP [20].
+Baselines.
+We compared the proposed method with four prior-based approaches which were developed on top of the IFM, including DCP [2], UDCP [3], GDCP [4] and DHL [8], two prior-based methods which did not consider the IFM, including Histogram Equalization (HE) and Retinex [11], and two supervised deep learning approaches, including WaterNet [16] and UColor [17].
+Data Sets.
+We conducted a series of experiments on five testing sets derived from four publicly available real-world underwater data sets: UIEB [16], SUIM-E [18], RUIE [43] and SQUID [8].
 The UIEB data set originally contains 890 pairs of degraded images, in which each reference image was selected from the results produced by applying 12 algorithms to a degraded image.
 Ten pairs of UIEB [16] images were discarded because the degraded images in these pairs had been included in the rest pairs.
-In total, we only used 880 pairs of UIEB images.
-The SUIM-E [31] data set comprises 1,525 pairs of degraded underwater images and a testing set of 110 labeled images.
-Three subsets are comprised of the RUIE [56] data set, including UIQS, UCCS and UHTS, which include 3,630, 300 and 300 unlabeled images, respectively.
-The SQUID [10] data set comprises 114 underwater images captured using a stereo camera, divided into the Michmoret, Katzaa, Nachsholim and Satil subsets.
-We utilized 2,185 images from the Underwater Scenes subset of the EUVP [20] data set.
-Two different PATS-UIENet models were trained for the real-world and synthetic testing images, referred to as $PATS\text{-}UIENet_{Real}$ and $PATS\text{-}UIENet_{Syn}$, respectively.
-Regarding both the models, we randomly selected 720 pairs from the 880 pairs of UIEB [22] images and 2,000 pairs from the Underwater Scenes subset of the EUVP [20] data set, respectively, which were used as the training images in the supervised learning scheme.
-We collected 1,934 unlabeled real-world underwater images, which cover diverse underwater scenes, and used these images in the unsupervised learning scheme for both the models.
+Only 880 pairs of UIEB images were used in this study.
+The SUIM-E data set comprises 1,525 pairs of degraded underwater images and a testing set of 110 labeled images.
+Three subsets are comprised of the RUIE data set, including UIQS, UCCS and UHTS, which include 3,630, 300 and 300 unlabeled images, respectively.
+The SQUID data set comprises 114 underwater images captured using a stereo camera, divided into the Michmoret, Katzaa, Nachsholim and Satil subsets.
+Regarding the supervised learning, we randomly selected 720 pairs from the 880 pairs UIEB [16] images and used these as the training set.
+For the unsupervised learning, we collected 1,934 unlabeled real-world underwater images, which cover diverse underwater scenes.
+In total, our training set contained 2,654 images.
 Among the rest of the UIEB [16] data set, 80 pairs of images were randomly selected as the validation set, while the remaining pairs were used as the testing set which is referred to as Test-U80.
-The UIEB data set also consists of 60 challenging images without references, which were used as the second testing set, referred to as Test-C60.
-A third testing set was obtained from the testing set of the SUIM-E data set [31], which contains 110 pairs of labeled images and is named as Test-S110.
+The UIEB data set also consists of 60 challenging images without reference, which were used as the second testing set, referred to as Test-C60.
+A third testing set was obtained from the testing set of the SUIM-E data set [18], which contains 110 pairs of labeled underwater images and is named as Test-S110.
 For the RUIE [43] data set, the UCCS was utilized as the fourth testing set, denoted as Test-UCCS.
-Fifty-three SQUID [10] images captured using the camera at the right hand side were randomly selected and were used as the fifth testing set, namely, Test-R53.
-The remaining 185 images in the Underwater Scenes subset of the EUVP [20] data set were used as the sixth testing set, referred to as Test-Scenes.
-IV-C Evaluation Metrics
-For the labeled testing data, we used Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [57] and Learned Perceptual Image Patch Similarity (LPIPS) [58] to assess the quality of an enhanced image with regard to the reference image.
-Regarding the unlabeled testing data, we used the Underwater Image Fidelity (UIF) [59] and Multi-scale Image Quality (MUSIQ) [60] metrics.
+Besides, 53 SQUID [8] images captured using the camera at the right hand side were randomly selected and were used as the fifth testing set, namely, Test-R53.
+Evaluation Metrics.
+For the labeled testing data, we used both Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [44] to assess the quality of an enhanced image with regard to the reference image.
+Regarding the unlabeled testing data, we used the Underwater Image Quality Measure (UIQM) [45] and Underwater Color Image Quality Evaluation (UCIQE) [46] metrics.
 When the SQUID [8] data set was tested, the Average Reproduction Angular Error ($\bar{\psi}$) [8] was used to evaluate the quality of color restoration.
 Since transmission maps are related to the scene distance, we used the Pearson Correlation Coefficient (PCC) calculated between the estimated transmission maps and the ground-truth depth map to assess the performance of transmission estimation [8].
-IV-D Implementation Details
-We implemented the proposed method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3090 graphics processing unit.
-During the training process, the images were resized to a resolution of $256\times 256$ pixels.
-The number of filters in the RCT was set to 16.
-For the $Enc\_1$ to $Dec\_1$ in the D-Stream and B-Stream, the numbers of filters were set to 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn.
+Implementation Details.
+We implemented our method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3080 graphics processing unit.
+During the training process, the images were resized to the resolution of $256\times 256$ pixels.
+The number of filters in the RCT was 16.
+For the $Enc_{1}$ to $Dec_{1}$ in the T-stream, the numbers of filters were 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn.
 Regarding the A-stream, the dimension of each token was 384 and the number of attention heads in the MHSA was 6.
-We trained the PATS-UIENet using the AdamW [61] optimizer.
-The learning rate and batch size were set to 1e-4 and 12, respectively.
-Four weighting factors, including $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{unsup}$, were set to 0.1, 0.005, 1 and 0.1, respectively.
-We first trained the PATS-UIENet using the bi-directional supervised scheme for 50 epochs as a warm-up stage.
-Then the semi-supervised scheme was used to train it for 1000 epochs.
-When the semi-supervised learning process was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for each epoch.
-V Experimental Results
-In this section, we will report the results obtained in the underwater image enhancement, transmission estimation and color restoration experiments and the ablation study.
-V-A Underwater Image Enhancement
-We evaluated the proposed PATS-UIENet along with sixteen baselines using three full-reference quantitative metrics and two non-reference quantitative metrics.
-A qualitative analysis and a performance analysis were also performed.
-In this subsection, the results obtained in the four experiments will be reported.
-V-A1 Full-Reference Quantitative Evaluation
-Since the Test-U80, Test-S110 and Test-Scenes testing sets contain reference images, a full-reference quantitative evaluation was conducted on the proposed method and the 16 baselines using the PSNR, SSIM [57] and LPIPS [58] metrics across these testing sets.
-The results are shown in Table I.
-As can be seen, our PATS-UIENet normally achieved the best performance across all three metrics, regardless of the testing set.
-Compared with the baselines designed on top of the IFM [1], such as UDCP [4] and DHL [8], and the existing semi-supervised learning approaches, including DDFormer [29], Semi-UIR [28] and UWFormer [30], our method normally showed a large margin.
-TABLE I: The full-reference quantitative evaluation of the proposed PATS-UIENet and sixteen baselines on three real-world testing sets.
-The best and second best results are highlighted in the Red Bold and Blue Italic fonts, respectively.
-Method Test-U80 Test-S110 Test-Scenes
+We trained the PA-UIENet using the AdamW [47] optimizer.
+The learning rate and batch size were set to 1e-4 and 6 respectively.
+Four weighting factors, including $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{unsup}$ were set to 0.001, 0.005, 10 and 0.001 respectively.
+We first trained the PA-UIENet using the bi-directional supervised scheme for 3,000 iterations as a warm-up stage.
+Then the semi-supervised scheme was used to train it for 150,000 iterations.
+When the semi-supervised learning was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for every 120 iterations.
+5 Experimental Results
+In this section, we will report the results obtained in the underwater image enhancement evaluation, parameter estimation and color restoration evaluation and ablation study.
+5.1 Evaluation on Underwater Image Enhancement
+The quantitative evaluation of eight baselines and the proposed method was conducted for the UIE task.
+The results are reported in Tables 1.
+As can be seen, our method outperformed both six prior-based methods and two supervised deep learning approaches on two labeled data sets, including Test-U80 and Test-S110, when the PSNR and SSIM metrics were used.
+On the other hand, the UIQM and UCIQE metrics were also used in the evaluation experiment where we did not utilize reference images.
+In this case, the three deep learning methods were inferior to the prior-based approaches.
+This finding is consistent with those reported in the existing studies [16, 18].
+Specifically, the performance of our method was normally comparable to that of the two deep learning approaches, i.e., WaterNet [16] and UColor [17], and was comparable to that of the prior-based methods.
+As demonstrated in [16, 18, 25], both the UIQM and UCIQE metrics are not always consistent with the quality of the images processed by UIE methods.
+It should be noted that, however, the prior-based methods did not produce stable results across the five testing sets.
+This should be attributed to the fact that these methods normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene is not suited to the prior assumption [4, 16].
+Table 1: The quantitative evaluation of eight baselines and our method on five real-world testing sets.
+The best and second best results are highlighted in the Red Bold and Blue Italic fonts respectively.
+Method Test-U80 Test-S110 Test-C60 Test-UCCS Test-R53
 […]
-UDCP [4] 9.51 33.66 41.74 10.07 34.29 37.66 14.76 56.43 30.31
-DHL [10] 15.16 63.93 30.40 14.61 62.93 28.39 14.99 69.73 22.36
-HE [9] 16.60 77.86 25.67 15.41 74.79 30.05 13.61 62.82 36.10
-HLRP [15] 13.56 22.76 33.90 12.55 29.24 33.18 12.12 18.21 39.35
-MMLE [14] 18.56 76.21 22.57 17.32 77.39 22.77 14.89 62.23 31.99
-WaterNet [22] 17.13 70.41 40.01 18.83 75.22 31.08 25.84 84.02 16.28
-UColor [23] 21.05 85.24 17.97 20.26 84.25 15.07 24.31 81.40 19.05
-FUnIE-GAN [20] 15.31 59.68 44.28 16.89 66.32 34.98 25.11 83.16 13.76
-USUIR [24] 16.92 68.69 28.63 17.54 76.58 21.38 15.14 66.51 32.44
-PUIE-Net (MP) [25] 22.26 88.89 12.01 21.85 89.19 9.70 21.85 73.47 24.45
-U-Transformer [21] 20.98 72.76 26.32 19.92 68.25 30.17 24.09 79.98 15.08
-URanker [26] 21.87 85.75 18.64 21.27 87.43 13.51 22.46 82.35 20.87
-CCMSRNet [27] 22.74 88.74 13.18 22.22 89.07 10.76 18.96 77.67 23.85
-DDFormer [29] 10.94 25.74 58.26 10.87 41.27 72.19 10.64 27.81 69.46
-Semi-UIR [28] 23.63 81.81 23.08 19.83 73.95 30.54 19.12 74.94 25.68
-UWFormer [30] 19.65 85.26 17.21 21.45 90.27 10.88 19.62 80.36 21.05
-PATS-UIENet (Ours) 23.59 90.16 10.42 22.78 90.54 8.85 25.87 84.34 12.04
-V-A2 Non-reference Quantitative Evaluation
-We used two non-reference metrics, i.e., UIF [59] and MUSIQ [60], to assess the performance of our method and sixteen baselines on the Test-U80, Test-S110, Test-Scenes, Test-C60, Test-UCCS and Test-R53 testing sets.
-The results are shown in Table II.
-It can be seen that our method achieved the best UIF score on three out of the six testing sets, including Test-U80, Test-S110 and Test-C60, and ranked the second on the Test-R53 testing set.
-Regarding the MUSIQ metric, the proposed method still demonstrated the competitive performance, especially on the challenging Test-S110, Test-Scenes and Test-R53 testing sets, even though it did not always produce the highest score.
-TABLE II: The non-reference quantitative evaluation of the proposed PATS-UIENet and sixteen baselines on six real-world testing sets.
-The best and second best results are highlighted in the Red Bold and Blue Italic fonts, respectively.
-Method Test-U80 Test-S110 Test-Scenes Test-C60 Test-UCCS Test-R53
-[…]
-UDCP [4] 36.40 44.99 32.01 51.38 61.25 36.99 25.41 33.03 34.72 29.72 3.24 36.85
-DHL [10] 41.32 47.43 14.80 57.22 0.55 36.57 25.04 37.27 3.62 31.07 2.53 42.64
-HE [9] 44.04 46.68 45.01 56.14 52.81 35.79 41.50 36.08 38.08 31.24 1.67 43.06
-HLRP [15] 0.59 49.73 0.91 56.87 0.71 40.21 3.57 34.65 0.32 34.44 1.26 43.53
-MMLE [14] 30.99 52.50 35.05 60.18 42.05 44.67 24.95 40.18 28.08 35.69 24.67 53.53
-WaterNet [22] 33.25 31.84 28.50 40.47 72.94 38.47 28.65 34.75 33.38 24.88 9.71 32.93
-UColor [23] 44.15 45.10 42.83 54.80 70.20 35.12 26.46 36.01 49.22 35.12 2.31 49.63
-FUnIE-GAN [20] 29.12 47.47 22.77 52.27 37.69 45.38 23.11 43.49 22.67 34.64 10.93 38.92
-USUIR [24] 5.17 44.44 37.02 56.83 48.82 37.91 22.35 34.15 20.54 28.83 17.98 39.44
-PUIE-Net (MP) [25] 9.18 49.05 60.98 60.44 27.69 45.46 3.39 38.69 51.76 29.18 46.16 45.15
-U-Transformer [21] 5.16 39.34 24.94 44.96 38.68 30.44 29.31 37.66 29.69 27.05 19.24 30.06
-URanker [26] 54.87 44.63 58.31 57.15 72.54 39.20 44.70 38.37 45.97 29.60 39.19 41.92
-CCMSRNet [27] 54.61 49.85 0.23 60.59 63.59 39.64 41.85 40.24 51.34 33.72 34.80 48.23
-DDFormer [29] -5.36 36.99 1.92 40.68 -3.16 23.87 11.14 36.21 -2.14 20.19 5.86 24.87
-Semi-UIR [28] 50.06 45.40 31.94 51.22 44.01 41.47 46.82 43.25 32.28 30.34 23.33 35.21
-UWFormer [30] 46.48 50.95 55.69 61.53 67.40 36.80 36.76 41.40 53.24 34.30 39.98 46.26
-PATS-UIENet (Ours) 58.83 49.91 61.20 60.66 68.07 46.70 46.99 40.49 51.18 31.63 42.51 46.63
-V-A3 Qualitative Analysis
-The enhanced images generated by the 16 baselines and our method on six testing sets are shown in Figs.
-4 - 9, respectively.
-It can be seen that some methods with the higher UIF [59] or MUSIQ [60] score did not produce visually pleasing results.
-For example, MMLE [14] produced the highest MUSIQ score on Test-U80 and Test-R53, but the resulting images suffered from pale colors and over-enhancement artifacts (see Figs.
-4 and 9).
-In contrast, our PATS-UIENet improved different degraded images and produced images with the natural and vivid color.
-Although the UIF [59] or MUSIQ [60] scores that our method produced were slightly lower than those obtained using some baseline methods [14, 22, 20] on certain testing sets, the enhanced images still manifest visually satisfactory quality.
-Figure 4: The results produced by 16 baselines and our method in terms of a degraded image in the Test-U80 testing set.
-Figure 5: The results produced by 16 baselines and our method in terms of a degraded image in the Test-S110 testing set.
-Figure 6: The results produced by 16 baselines and our method in terms of a degraded image in the Test-Scenes testing set.
-Figure 7: The results produced by 16 baselines and our method in terms of a degraded image in the Test-C60 testing set.
-Figure 8: The results produced by 16 baselines and our method in terms of a degraded image in the Test-UCCS testing set.
-Figure 9: The results produced by 16 baselines and our method in terms of a degraded image in the Test-R53 testing set.
-V-A4 Performance Analysis
-We compared the proposed PATS-UIENet with 11 deep baseline methods in terms of the number of parameters, Floating Point Operations Per Second (FLOPs) and inference speed (ms).
-Both the number of parameters and the FLOPs were calculated using the ptflops [^2] https://pypi.org/project/ptflops/ tool.
-As shown in Table III, the values of number of parameters, FLOPs and inference speed vary significantly.
-Specifically, UColor [23] exhibits the largest number of parameters (148.04M) and the largest FLOPs value (2804.36G), while UWFormer [30] incurs the slowest inference speed (491.00 ms).
-On the other hand, USUIR [24] has the lightest design with only 0.23M parameters and has a FLOPs value of 29.62G and owns the fastest inference speed (2.13ms).
-In contrast, our PATS-UIENet achieved a proper balance between the number of parameters or FLOPs and the inference speed.
-TABLE III: Comparison of between 11 deep UIE methods and our approach in terms of number of parameters, FLOPs and inference speed.
-Method #Params (M) FLOPs (G) Speed (ms)
-WaterNet [22] 1.09 142.84 6.40
-UColor [23] 148.04 2804.36 15.51
-FUnIE-GAN [20] 7.02 20.48 73.18
-USUIR [24] 0.23 29.62 2.13
-PUIE-Net (MP) [25] 1.40 70.54 101.89
-U-Transformer [21] 22.8 5.96 91.30
-URanker [26] 3.15 20.90 45.50
-CCMSRNet [27] 21.69 87.18 152.51
-DDFormer [29] 7.63 35.54 50.04
-Semi-UIR [28] 1.65 72.88 84.93
-UWFormer [30] 29.84 30.18 491.00
-PATS-UIENet (Ours) 45.71 179.70 86.17
-V-B Transmission Estimation and Color Restoration
-The quantitative evaluation of our methods and four prior-based baselines for transmission estimation and color restoration is reported in Table IV.
-It can be seen that our method achieved the superior, or at least the comparable, performance to that of the four prior-based baselines on four different subsets of Test-R53.
-Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets, no matter which metric was considered.
+DCP 12.48 63.34 1.35 0.51 14.52 67.38 1.36 0.57 0.56 0.51 0.03 0.44 -2.05 0.42
+UDCP 9.37 33.62 2.69 0.59 10.04 34.28 1.23 0.60 1.24 0.54 1.83 0.57 0.76 0.59
+GDCP 12.46 67.41 2.17 0.56 14.67 71.29 2.12 0.64 1.49 0.60 1.24 0.56 5.50 0.50
+DHL 14.49 62.81 3.19 0.64 14.96 69.68 2.35 0.66 1.85 0.64 2.98 0.63 0.22 0.64
+HE 16.48 77.77 3.58 0.67 15.37 74.76 2.97 0.68 2.32 0.65 3.93 0.67 1.71 0.64
+Retinex 1038 61.25 2.33 0.48 8.29 53.44 0.85 0.45 1.21 0.46 2.03 0.47 -2.03 0.51
+WaterNet 17.02 70.27 2.11 0.50 18.80 75.19 1.47 0.54 1.28 0.51 1.25 0.45 0.09 0.45
+UColor 20.93 85.19 3.44 0.60 20.23 84.23 2.30 0.63 2.08 0.58 3.22 0.55 1.28 0.54
+Ours 22.82 88.87 3.60 0.57 20.46 87.22 2.35 0.60 2.02 0.56 3.24 0.53 1.92 0.56
+The qualitative comparison of the results produced by the eight baselines and our method is shown in Fig.4.
+It can be seen that the enhanced images derived using our method presented the better visual quality, compared with the images enhanced by the baselines.
+Our method not only effectively removed the color casts in the degraded images but also greatly increased the visibility of them.
+In contrast, the prior-based methods usually introduced artifacts that decreased the visual quality while the learning-based methods could not eliminate the color casts completely.
+Figure 4: The results obtained using eight baselines and our method on five degraded images in Test-U80, Test-S110, Test-C60, Test-UCCS and Test-R53, respectively, shown from top to bottom.
+5.2 Evaluation on Transmission Estimation and Color Restoration
+The quantitative evaluation of our methods and eight baselines for transmission estimation and color restoration is reported in Table 5.2.
+It can be seen that our method achieved the superior, or at least the comparable, performance to that of the eight baselines on four different subsets of Test-R53.
+Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets no matter which metric was considered, while it performed comparably to the best baseline on the Nachsholim and Satil subsets.
 It should be noted that none of these methods produced a PCC value higher than 0.12 on the Satil subset.
@@
 As shown in Fig.
-10, our method was able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
-Figure 10: Each group shows a degraded image and the red channels of five transmission maps (top) and five enhanced images (bottom) obtained using four prior-based baselines and our method.
-TABLE IV: The quantitative evaluation of four prior-based baselines and our method on the four subsets of Test-R53, where PCC and $\bar{\psi}$ are used for transmission estimation and color restoration, respectively.
+5, our method is able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
+Table 2: The quantitative evaluation of eight baselines and our method on the four subsets of Test-R53, where PCC and $\bar{\psi}$ are used for transmission estimation and color restoration respectively.
 Method Michmoret Katzaa Nachsholim Satil
 […]
-DCP [3] -0.16 34.63 0.00 35.47 -0.26 34.75 0.03 36.21
-UDCP [4] -0.54 36.97 -0.15 40.43 0.06 38.80 0.12 51.45
-GDCP [5] 0.29 33.76 0.24 34.47 -0.04 34.28 -0.03 35.31
-DHL [10] -0.07 32.10 0.10 35.50 0.36 35.92 0.11 34.82
-Ours 0.62 12.83 0.32 9.47 -0.28 15.31 0.04 13.14
-V-C Ablation Study
-To investigate the effect of different components of the PATS-UIENet, we conducted a series of ablation experiments. For simplicity, we only utilized Test-U80 and the Michmorest subset in Test-R53.
-V-C1 Effect of the Semi-supervised Learning Framework
-To validate the effect of the proposed semi-supervised learning framework, we compared it with the unsupervised learning, supervised learning and bi-directional supervised learning frameworks.
-As reported in Table V, the proposed semi-supervised framework always outperformed the other frameworks in terms of different metrics across different testing sets.
-It has been suggested that our network can be better trained using the semi-supervised learning framework with both the labeled and unlabeled real-world images, which improves the enhancement performance.
-TABLE V: Comparison of different learning frameworks on Test-U80 and the Michmorest subset in Test-R53.
-Learning Framework Test-U80 Michmoret
+DCP -0.16 34.63 0.00 35.47 -0.26 34.75 0.03 36.21
+UDCP -0.54 36.97 -0.15 40.43 0.06 38.79 0.12 51.45
+GDCP 0.29 33.76 0.24 34.47 -0.04 34.28 -0.03 35.31
+DHL -0.07 32.10 0.10 35.50 0.39 35.91 0.11 34.82
+HE - 35.44 - 35.43 - 34.37 - 36.33
+Retinex - 31.44 - 34.37 - 33.29 - 35.31
+WaterNet - 25.40 - 25.70 - 25.80 - 28.31
+UColor - 15.55 - 13.94 - 17.21 - 21.43
+Ours 0.71 11.51 0.48 8.75 0.18 19.59 0.04 13.46
+Table 3: The ablation study on Test-U80 and the Michmorest subset in Test-R53.
+Method Test-U80 Michmoret
 […]
-Unsupervised 16.85 78.44 29.91 0.50 5.46
-Supervised 22.19 88.23 13.07 0.57 17.01
-Bi-supervised 23.00 89.49 11.93 0.61 15.02
-Semi-supervised (Ours) 23.59 90.16 10.42 0.62 12.83
-V-C2 Effect of the RCT and RCM
-To examine the effect of the RCT and RCM on the performance of our PATS-UIENet, we removed the RCT, the RCM or both of them.
-In this case, we derived three variants of the PATS-UIENet: one with both the RCT and RCM removed, one with only the RCT removed and one with only the RCM removed, denoted as “Simplified”, “w/o RCT” and “w/o RCM”, respectively.
-As shown in Table VI, the complete PATS-UIENet achieved the best performance in terms of all the metrics.
-Removal of the RCT led to a slight increase in the PSNR and SSIM metrics compared to the Simplified version, but worsened the LPIPS and $\bar{\psi}$ scores.
-On the other hand, removal of the RCM reduced the performance with regard to all metrics across the two data sets, confirming its usefulness in facilitating information exchange between the B-Stream and A-Stream.
-The results demonstrate that both the RCT and RCM contribute meaningfully to the performance of our method.
-TABLE VI: Impact of the RCT and RCM on the performance of our method on Test-U80 and the Michmorest subset in Test-R53.
-Variant Test-U80 Michmoret
-[…]
-Simplified 22.96 89.64 11.25 0.59 13.59
-w/o RCM 22.10 88.83 12.20 0.60 13.50
-w/o RCT 23.14 90.03 12.20 0.60 13.97
-Ours 23.59 90.16 10.42 0.62 12.83
-V-C3 Effect of the Degradation Control Factor
-To evaluate the impact of the degradation control factor $\alpha$ on the performance of the PATS-UIENet, we conducted an ablation experiment on testing three different values of $\alpha$.
-As shown in Table VII, the default setting of $\alpha$ = 0.1 achieved the best performance with regard to both the full-reference and the non-reference metrics across the two data sets.
-TABLE VII: The effect of the degradation control factor, $\alpha$, on the performance of our method.
-$\alpha$ Test-U80 Michmoret
-[…]
-0.001 23.21 89.77 11.02 0.59 13.92
-1 23.06 89.72 11.08 0.61 14.11
-0.1 (Ours) 23.59 90.16 10.42 0.62 12.83
-V-C4 Effect of Loss Hyperparameters
-To evaluate the impact of the hyperparameters used for the loss functions, including $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{unsup}$, on the performance of the proposed PATS-UIENet, we conducted a comprehensive ablation study by changing their values.
-Regarding the supervised learning loss function (see Eq.
-(7)), the results produced by our method with different combinations of the $\lambda_{1}$ and $\lambda_{2}$ values are shown in Table VIII.
-It can be seen that the combination that we used produced the best result in terms of each metric.
-For the unsupervised learning loss function (see Eq.
-(15)), the results produced by our method with different values of $\lambda_{3}$ are reported in Table IX.
-Again, the value that we chose led to the best result no matter what metric was considered.
-In terms of the semi-supervised learning loss function (see Eq.
-(16)), we present the results obtained using our method with different $\lambda_{unsup}$ values in Table X.
-As can be observed, the value that we utilized produced the best result with regard to each metric.
-The above findings highlight the effectiveness of the values of different loss hyperparameters that we chose.
-TABLE VIII: The effect of different combinations of hyperparameters used for the loss function of the supervised learning scheme, i.e., $\lambda_{1}$ and $\lambda_{2}$, on the performance of our method.
-$\lambda_{1}$ $\lambda_{2}$ Test-U80 Michmoret
-[…]
-1 0.0005 22.04 88.01 12.94 0.36 16.92
-0.0001 0.0005 22.84 89.65 10.85 0.61 13.57
-0.1 0.1 22.95 89.67 11.52 0.61 15.69
-0.1 0 22.91 80.57 11.54 0.59 16.92
-0.1 0.0005 (Ours) 23.59 90.16 10.42 0.62 12.83
-TABLE IX: The effect of the weighting factor used for the loss function of the unsupervised learning scheme, i.e., $\lambda_{3}$, on the performance of our method.
-$\lambda_{3}$ Test-U80 Michmoret
-[…]
-0.1 22.95 89.80 11.01 0.61 15.14
-10 22.36 88.19 13.09 0.60 14.72
-1 (Ours) 23.59 90.16 10.42 0.62 12.83
-TABLE X: The effect of the weighting factor used for the loss function of the semi-supervised learning scheme, i.e., $\lambda_{unsup}$, on the performance of our method.
-$\lambda_{unsup}$ Test-U80 Michmoret
-[…]
-1 22.32 88.75 12.16 0.61 14.08
-0.01 23.31 89.96 10.92 0.61 12.84
-0.1 (Ours) 23.59 90.16 10.42 0.62 12.83
-V-C5 Effect of Different Stream Architectures
-To further examine the effect of different stream architectures on the PATS-UIENet, we constructed its three variants by building the D-Stream, B-Stream and A-Stream using convolution and/or Transformer networks.
-In addition, we obtained a fourth variant by removing the D-Stream.
-In essence, this variant is equal to the network built on top of the original IFM [1].
-The four variants were trained and tested using the same setup as that used for our PATS-UIENet.
-The results produced by these variants and our method are shown in Table XI.
-It can be seen that our method produced the better, or at least comparable, results in contrast to the four variants.
-This finding supports the design philosophy of the PATS-UIENet.
-That is to say, CNNs are suitable for capturing local characteristics while Transformer is good at encoding global characteristics.
-Moreover, our network which was built on top of the revised IFM [2] normally outperformed the variant constructed based on the original IFM [1], confirming the choice of the revised IFM [2].
-TABLE XI: Comparison between our method and its four variants obtained by building the D-Stream, B-Stream and A-Stream using different networks.
-Method Streams Test-U80 Michmoret
-[…]
-$\text{Variant}_{1}$ CNN CNN CNN 22.31 88.69 12.48 0.58 12.44
-$\text{Variant}_{2}$ Trans Trans CNN 20.71 75.36 25.31 0.73 17.28
-$\text{Variant}_{3}$ Trans Trans Trans 20.57 75.00 25.73 0.71 14.23
-$\text{Variant}_{4}$ N/A CNN Trans 22.51 88.85 11.47 0.71 15.95
-Ours CNN CNN Trans 23.59 90.16 10.42 0.62 12.83
-VI Conclusion
-In this paper, we introduced a novel Physics-Aware Triple-Stream Underwater Image Enhancement Network, namely, PATS-UIENet, which explicitly estimates the degradation parameters of the revised IFM, for the UIE task.
+Simplified 22.16 89.57 -0.15
+w/ RCT 22.97 89.93 0.58
+w/ Bi-Sup 22.51 88.94 0.68
+w/ Semi-Sup 22.82 88.87 0.71
+Figure 5: Each group shows a degraded image and the red channels of five transmission maps (top) and five enhanced images (bottom) obtained using four prior-based baselines and our method.
+5.3 Ablation Study
+To examine the effect of each component of our method, including the RCT, bi-directional supervised scheme and semi-supervised learning framework, we conducted an ablation study.
+For simplicity, we only utilized Test-U80 and the Michmoret subset in Test-R53.
+The PA-UIENet without the RCT together with the forward-enhancement supervised learning module was crafted as a simplified version.
+On top this version, we derived three additional variants by inserting the RCT (w/ RCT) and then replacing the forward-enhancement supervised module by the bi-directional supervised scheme (w/ Bi-Sup) or the semi-supervised learning framework (w/ Semi-Sup).
+As shown in Table 5.2, the simplified version outperformed all the baselines in terms of the PSNR and SSIM metrics.
+However, this variant did not estimate transmission maps well, which may be attributed to the overfitting problem resulted from limited training data.
+When the RCT was used, the performance of image enhancement and transmission estimation was boosted.
+It was indicated that the adaptability of the red channel increased the robustness and generalization of the model.
+Though the utilization of the bi-directional-supervised learning scheme or the semi-supervised learning framework had a negative effect on the PSNR and SSIM values, the effectiveness of transmission estimation continuously increased.
+It is suggested that the use of the backward-degradation learning module or the unsupervised learning scheme alleviates the overfitting problem.
+6 Conclusion
+In this paper, we developed a novel Physics-Aware Dual-Stream Underwater Image Enhancement Network, namely, PA-UIENet, which explicitly estimates the degradation parameters of the IFM, for the UIE task.
 To overcome the challenge of insufficient data, we also adopted an IFM-inspired semi-supervised learning framework, comprising a bi-directional supervised scheme and an unsupervised scheme.
-Due to the complementary action of the two schemes, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model is stronger, compared to supervised and unsupervised methods.
-To our knowledge, this study made the first effort to jointly exploit the physics-aware deep network and the IFM-inspired semi-supervised learning technique for the UIE task.
-The proposed method performed better than, or at least comparably to, sixteen baselines on six underwater testing sets in different evaluations.
+Due to the complementary action of the two schemes, the PA-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model is stronger, compared to supervised and unsupervised methods.
+To our knowledge, this study made the first attempt to jointly exploit the physics-aware deep network and the IFM-inspired semi-supervised learning technique for the UIE task.
+The proposed method performed better than, or at least comparably to, eight baselines on five real-world underwater testing sets in different evaluations.
 The promising results should be due to the fact that our method is able to not only estimate degradation parameters but also learn the characteristics of diverse underwater scenes.